### v2

# Know You First and Be You Better: Modeling Human-Like User Simulators via Implicit Profiles

Abstract
User simulators are crucial for replicating human interactions with dialogue systems, supporting both collaborative training and automatic evaluation, especially for large language models (LLMs). However, existing simulators often rely solely on text utterances, missing implicit user traits such as personality, speaking style, and goals. In contrast, persona-based methods lack generalizability, as they depend on predefined profiles of famous individuals or archetypes. To address these challenges, we propose User Simulator with implicit Profiles (USP), a framework that infers implicit user profiles from human-machine conversations and uses them to generate more personalized and realistic dialogues. We first develop an LLM-driven extractor with a comprehensive profile schema. Then, we refine the simulation through conditional supervised fine-tuning and reinforcement learning with cycle consistency, optimizing it at both the utterance and conversation levels. Finally, we adopt a diverse profile sampler to capture the distribution of real-world user profiles. Experimental results demonstrate that USP outperforms strong baselines in terms of authenticity and diversity while achieving comparable performance in consistency. Furthermore, dynamic multi-turn evaluations based on USP strongly align with mainstream benchmarks, demonstrating its effectiveness in real-world applications  [^1] All resources will be released at https://github.com/wangkevin02/USP.
{CJK}
UTF8gbsn
Know You First and Be You Better:
Modeling Human-Like User Simulators via Implicit Profiles
Kuang Wang1,2, Xianfei Li1, Shenghao Yang1, Li Zhou1, Feng Jiang1,2,3 thanks: Feng Jiang is the corresponding author., Haizhou Li1,2 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 3Shenzhen University of Advanced Technology kuangwang@link.cuhk.edu.cn, jeffreyjiang@cuhk.edu.cn
1 Introduction
Figure 1: Examples of performance across various user simulators in multi-turn human-LLM interactions. OF and SC refer to objective facts and subjective characteristics, respectively.
The user simulator is designed as a proxy for real users in interactions with large language models (LLMs). It can simulate realistic user by generating the target user’s behavior or utterances based on the specified characteristics, enabling dynamic multi-turn interactions with LLMs Wan et al. (2022) and scene reproduction Wang et al. (2024b). As a result, it becomes an effective alternative Liu et al. (2023); Ferreira et al. (2024) in scenarios where real-world human-computer interaction data is difficult to obtain, especially in domains with privacy and ethical concerns, such as medical consultations Valizadeh and Parde (2022). It also helps Simulation-to-Reality (Sim2Real) applications, such as tutorial strategies, election simulations, and public opinion research Liu et al. (2024); Zhang et al. (2024b); Chuang et al. (2024).
Recent LLMs advance promote user simulators by enhancing naturalness and utility Deng et al. (2024); Zhang et al. (2024a), as shown in Figure 1. Since directly using LLM as the user simulator suffers from role confusion Xu et al. (2023a), some works Xu et al. (2023a); Kong et al. (2024); Sun et al. (2024) attempt to enhance its authenticity by training a user simulator on the conversation dataset. However, they are only trained on text utterance, making it difficult to simulate diverse user behaviors without seed context, exhibiting limited self-awareness Tseng et al. (2024) and failing to maintain a consistent personality. Although some LLM-based role-playing methods Moon et al. (2024) utilize predefined profiles to alleviate this problem, they require additional extensive annotations and can only be applied to celebrities, lacking the diversity of simulation.
To address the issues above, we believe that a user simulator knows users’ intrinsic characters hidden in their conversations first and then can be a better simulation. Therefore, we treat user simulation as a dialogue reconstruction task and propose a novel framework named the User Simulator with implicit Profile (USP). It is decomposed into implicit profile extraction to capture the user’s underlying characteristics from the target user dialogue and conditional generation based on the profile.
In this framework, we first propose an LLM-driven profile extractor to extract implicit profiles from user conversations with a well-designed profile schema. Inspired by interpersonal interaction theoryKruglanski and Higgins (2013), our profile schema contains two dimensions (objective facts (OF) and subjective characteristics (SC)) with a dozen attributes to describe the user comprehensively. Different from existing works Cheng et al. (2024b); Tu et al. (2024), we then polish the profile attributes into natural, descriptive profiles to ensure generalization.
Then, we integrate the extracted user profiles into the user simulator through two-stage training: (1) conditional supervised fine-tuning with user profiles for utterance-level simulation, and (2) reinforcement learning with cycle consistency to align reflected profiles from simulated dialogues with given profiles for conversation-level simulation. We also implement a diverse profile sampler to capture authentic user distributions.
Our experiments reveal that USP enhances semantic and stylistic similarity of reconstructed multi-turn dialogues to target dialogues by about 34% and 43% over the leading baseline, with reconstruction errors halved, showcasing superior authenticity and diversity. It matches the consistency of ProfileGPT(4o), a GPT-4o-based simulation baseline, boosting dialogue profile consistency by 14% in multi-turn scenarios while performing comparably in single-turn cases. Furthermore, our USP-based multi-turn dynamic evaluation of LLMs for downstream tasks aligns well with established benchmarks,allowing for a more granular assessment of LLM performance across different user groups. Our key contributions are outlined below:
  • We propose a novel approach for constructing user simulators using implicit user profiles embedded in human-LLM conversations.
  • We develop a new framework that infers implicit user profiles as insight, further enhanced with conditional fine-tuning and reinforcement learning with cycle consistency for better simulation at both the utterance and conversation levels.
  • Experimental results show USP surpasses all baselines in authenticity and diversity, maintains comparable consistency, and proves effective for multi-turn dynamic evaluation of LLMs.
2 Related Works
2.1 General User Simulator
Early user simulators including agenda-based methods Schatzmann et al. (2007); Schatzmann and Young (2009) and model-based methods  Asri et al. (2016); Kreyssig et al. (2018). These simulators were initially designed with a narrow scope due to limited natural language generation capabilities, such as generating synthetic binary preference responses Christakopoulou et al. (2016) in conversational recommendation systems.
Recent advancements in LLMs enabled more sophisticated simulations of realistic conversations, offering significantly enhanced natural language flexibility. These advances include the use of LLMs for self-chat Xu et al. (2023b) and dual LLM architectures, where separate models role-play user and assistant based on seed conversations Ding et al. (2023). Following these innovations, other trained user simulators, such as PlatoLM Kong et al. (2024) and Parrot Sun et al. (2024), learn human discourse patterns directly from human-LLM interactions in conversations.
2.2 Persona-based User Simulator
General user simulators often struggle to capture the full spectrum of diverse user needs, leading to a growing interest in persona-based personalization to improve both controllability and diversity in simulations Takanobu et al. (2020). Some researchers attempt to leverage goal generators Takanobu et al. (2020) to create diverse user goals or retrieval-based personas derived from historical data Shi et al. (2019) to guide user simulators in task-oriented dialogue (ToD) systems.
With the rise of LLMs and their impressive zero-shot role-playing abilities Njifenjou et al. (2024), prompt-driven user simulation has become the dominant approach. For example, LLMs have been used with carefully designed predefined profiles to align with human beliefs Chuang et al. (2024), simulate consultation scenarios with users exhibiting varying personalities and needs in ToD systems Zhang et al. (2024a), and model user preferences in conversational recommendation systems Yoon et al. (2024).
3 Task Definition
We formulate user simulation as a dialogue refactoring task, aiming to replicate multi-turn user behavior in target dialogues. Given a target dialogue $d_{i}=\{(u_{i1},r_{i1}),\dots,(u_{ij},r_{ij})\}$ between a user $U_{i}$ and a response model $R_{i}$, where $u_{ij}$ and $r_{ij}$ represent the $j$-th turn user utterance and the corresponding model response, respectively.
To achieve high-fidelity simulation of user responses within a given context, we aim to minimize the utterance-level distance $D_{\text{utt}}(u_{ij},u_{ij}^{\prime})$ where $u_{ij}^{\prime}\sim P_{prob}(\cdot|c_{ij}^{\prime},U_{i}^{\prime})$. Here, $c_{ij}^{\prime}$ represents the context used by the user simulator $U_{i}^{\prime}$ to generate $u_{ij}^{\prime}$. This optimization ultimately leads to reducing the dialogue-level distance, as formulated in Eq. 1.
$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime})}D_{\text{dia}}(d_{i},d_{i}^{\prime})$ (1)
where $D_{\text{dia}}$ represents a distance function that evaluates the user utterances between the simulated dialogue $d_{i}^{\prime}$ and the real conversation $d_{i}$.
Recent studies show that role-playing with specific user profiles ($P_{i}$) can effectively achieve diverse user simulations Liu et al. (2023). However, unlike celebrities or well-known characters, user profiles in real-world conversations are often implicit and difficult to obtain Wang et al. (2024a).
To address this, we reformulate the task by first extracting the implicit user profile from the given dialogue using profile extractor $P_{\text{extractor}}$, and then reconstructing a closer dialogue, with an emphasis on the user’s utterances, as described in Eq. 2.
$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime},P_{i})}D_{\text{dia}}(d_{i},d_{i}^{\prime}),$ (2)
where $P_{i}=P_{\text{extractor}}(d_{i})$.
4 Modeling User Simulator with Implicit Profiles
Figure 2: Overview of our proposed User Simulator with implicit Profile(USP) framework.
To accomplish this task, we propose the User Simulator with Implicit Profiles (USP) framework, as illustrated in Figure 2, which aims to minimize the objective in Eq. 2 while ensuring authenticity, consistency, and diversity.
4.1 User Profile Construction
Category Dimension Attributes
Objective Facts
Scene-Consistent
Attributes
Age, Gender,
Location, Occupation,
Education, Family Relationship,
Routines/Habits, Social
Relationships, Other Experiences
Scene-Related
Attributes
Goals/Plans, Task Details
Subjective
Characteristics
Intrinsic
Characteristics
Big Five Personality Traits,
Language Styles
Table 1: The Designed User Profile Schema.
4.1.1 User Profile Schema
We believe that the user profile should reveal user characteristics from two aspects: explicit personal information and implicit communication styles. Therefore, inspired by interpersonal interaction theory Zhou et al. (2024b), we design a user profile schema containing objective facts(OF) and subjective characteristics(SC) to represent them, as shown in Table 1.
The OF focuses on common topics in human conversation (Cheng et al., 2024b; Dunbar et al., 1997) including Scene-Consistent Attributes (such as age, gender, and location) and Scene-Related Attributes (such as goal,and task details). SC considers both external and internal personality dimensions represented by language style Wang et al. (2024a) and the Big-Five Traits Different from previous work Cheng et al. (2024b); Tu et al. (2024), we reformulate discrete attributes into coherent narrative descriptions to achieve greater generalization and flexibility.
4.1.2 User Profile Extractor
To obtain such a user profile, we design an LLM-driven user profile extractor extracting the implicit user profile from the human-LLM conversation. The extractor first leverages advanced LLM (such as GPT-4o) to extract the user character attributes mentioned above with a well-designed prompt. Then, the extractor collects the valid attributes (No empty) together and polishes them into natural language descriptions. Further prompt details regarding the extractor can be found in Appendix A.2.
4.1.3 Profile Quality Verification
Due to the reliance on predefined user profiles in existing role-playing approaches Zhou et al. (2024b), the correlation between user profiles and conversations has been largely overlooked. To address this limitation, we propose Dialogue Profile Consistency (DPC) for evaluating OF based on atomic fact verification that measures both precision (DP.P) and recall (DP.RJandaghi et al. (2024).
Drawing inspiration from FactScore Min et al. (2023) and Con.Score Madotto et al. (2019), we first introduce Factual Consistency (Fact.Con), as defined in Eq. 3. Given a target $T$, we evaluate the consistency between the source and target by decomposing $T$ into atomic facts $af_{k}$ using an atomic fact generator (afg). We then compute the natural language inference (NLI) score for each atomic fact with respect to the source $S$.
[…]
Additionally, we use a validation score (Val.Score) to assess the quality of SC descriptions based on the dialogue, rating them on a scale from 1 to 5 using GPT-4o. Detailed prompts in Appendix D.
4.2 Conditional Supervised Fine-Tuning
To empower the LLM with the general capability to simulate diverse users at the utterance level, we utilize conditional supervised fine-tuning based on user profiles. It enables the LLM to learn the conditional generation mapping based on both the extracted profile $P_{i}$ and context $c_{ij}^{\prime}$. As subtle misalignment between the core objectives of the user simulator and the response model, the SFT language modeling loss is adjusted as shown in Eq. 4.
$l_{\text{lm}}=\sum_{j}\sum_{k}-\log P_{prob}(u_{i,j,k}^{\prime}|u_{i,j,<k}^{\prime},c_{i,j}^{\prime},P_{i})$ (4)
where $u_{ijk}$ represents the $k$-th token of the $j$-th utterance from the $i$-th user.
4.3 Reinforcement Learning with Cycle Consistency
To further enhance conversation-level consistency, we introduce Reinforcement Learning with Cycle Consistency (RLCC), which optimizes the user simulator by aligning the reflected profile, extracted from simulated dialogues, with the target profile.
In this stage, we extract the simulator profile $P_{i}^{\prime}$ from dialogues $D^{\prime}$ generated by the user simulator $U_{i}^{\prime}$ using the profile generator, based on virtual profiles sampled by our diverse profile sampler (see Section 4.4). Our goal is to maximize the semantic similarity between the target profile $P_{i}$ and the extracted reflected profile both in objective facts and subjective characteristics. The dialogue-level reward is then distributed to each user utterance within the dialogue, as shown in Eq. 5.
$r^{cc}_{i,j}=\text{sim}(P_{i},P^{\prime}_{i})$ (5)
Then, we optimize profile recall through Proximal Policy Optimization (PPO) Schulman et al. (2017) using cycle consistency as a dialogue-level reward signal. It enhances the user simulator’s self-expression in dialogues, moving beyond simple profile adherence or precision consistency.
Additionally, to prevent reward hacking, we incorporate an AI detection model as an auxiliary reward. The final reward is formulated in Eq. 6.
$r_{i,j}=\lambda r^{cc}_{i,j}+(1-\lambda)r^{ai\_detect}_{i,j}$ (6)
where $r^{ai\_detect}_{i,j}=\text{AI\_detect}(u_{i,j})$ and $\lambda=0.8$ is used to emphasize the importance of cycle consistency. The AI detection model Yang et al. (2024) and profile generator are all fine-tuned based on our train dataset, details in Appendix B.1.
4.4 Diverse Profile Sampling
To generate diverse and naturalistic user profiles that reflect real-world distributions, we propose a Density Profile Sampler that preserves distribution characteristics while ensuring coverage of underrepresented cases. Profiles are first embedded into the semantic space using SimCSE Gao et al. (2021), then projected with UMAP McInnes et al. (2018) to maintain density relationships. We then estimate the underlying profile distribution using Gaussian Kernel Density Estimation (GKDE). This enables direct sampling of realistic profiles based on probability density, while generating virtual profiles during random sampling by combining OF and SC descriptions from nearest neighbors.
5 Experiments
We evaluate user simulators’ authenticity and consistency at both the utterance and conversation levels while assessing diversity through the difference between our simulated and real user distributions.
5.1 Datasets
We select the popular LMSYS-Chat-1M Zheng et al. (2023) as our data source for the experiment, which contains one million human-LLM conversations. Following previous work Kong et al. (2024), we filter the samples for non-English language, toxicity, and redundancy and obtain a complete 94,874 samples (87,882/4,626/2,366 for Training/Validation/Test datasets). Then, we use a GPT-4o-based profile extractor described in Section 4.1 to preprocess them with annotating user profiles for each conversation and construct them into LMSYS-USP. Detailed preprocessing are in Appendix A.1.
Dataset Profile Source OF SC
[…]
LMSYS-USP GPT4o 86.89 25.64 82.24 3.71 84.50 4.42
LMSYS-USP Distill-llama3 86.15 23.81 81.95 3.71 84.00 4.36
Persona Chat GPT4o 86.21 22.82 62.76 7.86 72.64 4.35
Persona Chat Human 76.21 8.59 42.94 7.86 54.93 -
ConvAI2 GPT4o 68.71 17.44 39.15 9.97 49.88 3.47
ConvAI2 Human 25.69 8.70 12.64 9.97 16.94 -
Table 2: Automatic evaluation results of profile quality across different datasets.
We used DPC and Val.Score to automatically evaluate the quality of extracted user profiles on the test set of LMSYS-USP, plus Persona-Chat Zhang et al. (2018) and ConvAI2  [^2] We use the human-to-bot dataset in https://huggingface.co/datasets/convai-challenge/conv_ai_2 Dinan et al. (2019) with manually annotated dialogue datasets of other existing predefined profiles. Table 2 shows that the extracted profile can achieve over 84% DPC and even the distill-llama3 is close to that of GPT-4o, demonstrating the effectiveness of our annotation. Additionally, we select 100 samples for manual evaluation and it shows over 4/5 scores for the quality of the generated profiles (See Appendix B.4 for further details).
5.2 Baseline Models
(1) User Simulator without User Profile: This includes the untrained DialogueGPT(4o), where GPT-4o relies solely on context to predict the next user utterance, and PlatoLM Kong et al. (2024), which is fine-tuned on our training dataset using LLaMA-3-8B and can be considered equivalent to our approach without incorporating profile.
(2) User Simulator Guided by the User Profile: We adopt ProfileGPT(4o) and ProfileGPT(llama), which utilizes GPT-4o and LLAMA-3-8B-Instruct AI@Meta (2024) as the role-play backbone with our constructed profiles. Additionally, we include CharacterGLM Zhou et al. (2024a), which performs role-playing with any given profile, and CharacterLLM Shao et al. (2023), which role-played on several famous celebrities. The detailed setup and prompts are provided in Appendix B.3.
5.3 Metrics
Authenticity: We use SimCSE Gao et al. (2021) to compute semantic similarity (Sem-Sim) and style embeddings Wegmann et al. (2022) to compute style similarity (Style-Sim) for evaluating $D_{\text{utt}}(u_{ij},u^{\prime}_{ij})$ and $D_{\text{dia}}(d_{i},d_{i}^{\prime})$. We also employ Author Verification Accuracy (AVA) to assess stylistic consistency by measuring whether paired sentences share authorship based on similarity thresholds Wegmann et al. (2022). For multi-turn evaluation, we compute dialogue-level distances by concatenating each user’s utterances.
Consistency: We evaluate dialogue-level consistency between dialogues and profiles using reverse metrics: r-DP.P and r-DP.R, calculated as $Fact.Con(P_{i},d_{i}^{\prime})$ and $Fact.Con(d_{i}^{\prime},P_{i})$ respectively, along with their harmonic mean r-DPC, from a profile-centric perspective. For utterance-level evaluation, we compute the mean DP.P score across individual utterances. Additionally, we incorporate Persona Coverage (P.Cover) Song et al. (2019) to assess keyword-level consistency. The Subjective Characteristic Score (SC.Score), measures the reflection of subjective traits, assessed by GPT-4o with prompt in Appendix D.
Diversity: We compute the Absolute Difference Value (ADV), which represents the Euclidean distance between the PCA-reduced embeddings of generated and target dialogues, to evaluate the discrepancy between the distribution of the reconstructed dialogues and the original dialogues.
Additionally, we assess multi-turn dialogue continuity using the early stop rate (ESR), which flags premature endings caused by repetitive responses or repeated gratitude expressions over three turns.
5.4 Results
5.4.1 Utterance-Level Evaluation
Model Type Model Authenticity Consistency
[…]
w/o Profile DialogueGPT(4o) 40.24 13.75 11.28
PlatoLM 39.37 43.11 40.29
With Profile Character_LLM 37.54 18.88 15.03 54.77 66.62 2.43
Character_GLM 38.51 22.28 18.17 68.72 57.72 2.95
ProfileGPT(llama) 39.82 14.88 13.47 82.19 72.29 3.92
ProfileGPT(4o) 41.66 5.74 9.87 92.73 73.34 4.71
USP w/o RLCC 54.25 46.57 43.61 71.30 71.56 3.36
USP 53.38 46.60 43.35 72.61 71.23 3.39
Table 3: Utterance-level performance comparison of different models on authenticity and consistency metrics.
Model Type Model Continuity Authenticity Consistency
[…]
w/o Profile DialogueGPT(4o) 35 48.91 14.21 10.58
PlatoLM 18 43.24 32.43 31.60
With Profile Character_LLM 52 23.37 7.13 4.69 25.48 6.43 10.27 21.49 2.82
Character_GLM 44 40.19 10.86 12.67 39.51 29.61 33.85 42.75 3.64
ProfileGPT(llama) 31 46.84 10.58 11.63 67.09 29.98 41.44 47.72 4.19
ProfileGPT(4o) 32 48.87 10.15 11.26 76.59 43.72 55.66 51.02 4.56
USP w/o RLCC 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
USP 10 65.39 46.23 38.77 56.24 74.38 64.05 44.08 3.35
Table 4: Conversation-level performance comparison of different models on authenticity and consistency metrics.
In the utterance-level evaluation, we evaluate the quality of a single-turn response generated by the testing models in a given context.
As shown in Table  3, USP outperforms all baselines in terms of authenticity, with 53.38 and 46.60, as measured by both semantic (Sem-Sim) and stylistic (Style-Sim) similarity metrics. This highlights the effectiveness of our implicit profile-assisted approach for user-LLM dialogue reconstruction, particularly when compared to context-only models like PlatoLM. Although dedicated role-playing models (ProfileGPT variants) achieve higher consistency scores (r-DP.P), this can be attributed to their direct profile text copying. USP maintains comparable overall performance while striking a better balance between authenticity and consistency.
5.4.2 Conversation-Level Evaluation
In the conversation-level evaluation, we assess the performance of testing models to chat with GPT-4o in multi-turns, providing the profile or the first turn of the reference dialogue according to their needs.
As shown in Table 4, USP outperforms baseline models in authenticity, consistency, and continuity. With the lowest ESR(10), it demonstrates superior dialogue continuity. Notably, USP’s advantage in authenticity is more pronounced in conversation-level scenarios than sentence-level predictions. In terms of consistency, USP demonstrates exceptional performance in r-DP.R metrics and achieves significantly higher r-DPC scores for overall profile dialogue consistency. This superior performance, particularly when compared to role-playing models such as ProfileGPT(4o) and ProfileGPT(llama) which show high P.Cover, suggests that RLCC effectively captures abstract profile characteristics beyond merely matching surface-level keywords. See Appendix C.2 for further case study analysis.
5.4.3 Human Evaluation
We also randomly selected 100 samples with 8 evaluators to conduct conversation-level human evaluations considering authenticity and consistency. Authenticity was assessed through Style, Semantics, and Quality, while consistency covered Accuracy, Completeness, and Quality. Full evaluation details are provided in Appendix B.4.
Table 5 demonstrates USP’s superior performance in both authenticity and consistency metrics. Our USP significantly outperforms ProfileGPT (4o) in terms of authenticity (74 vs. 13) and consistency (61 vs. 35) in manual evaluation. USP is superior to PlatoLM trained on the same dataset in terms of authenticity, which demonstrates the usefulness of implicit profile modeling. Thanks to the RLCC module, our USP model has significantly improved consistency (43 vs. 30) by aligning user profiles.
Baseline Metrics (USP win/tie/loss)
Authenticity Consistency
($\kappa$=0.548) ($\kappa$=0.561)
ProfileGPT(4o) 74/13/13 61/4/35
PlatoLM 55/12/33 -
USP w/o RLCC 37/32/31 43/27/30
Table 5: Human evaluation results comparing baselines with USP on authenticity and consistency.
5.4.4 Diversity Sampling Evaluation
Model Configuration Continuity Authenticity Consistency
[…]
USP w/o RLCC 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
USP (5:5) 14 66.28 41.22 37.03 52.23 71.59 60.39 43.58 3.55
USP (8:2) 10 65.39 46.23 38.77 56.24 74.38 64.05 44.08 3.35
USP (9:1) 12 66.91 38.87 33.62 58.36 70.62 63.90 46.75 3.33
Table 6: Ablation study of hyperparameters in RLCC.
Figure 3 shows the absolute difference value between target dialogue and generated dialogue by various models across different percentiles. From the results, we observe that USP and USP w/o RLCC consistently achieve the smallest ADV across all percentiles, indicating the dialogues they generated that closely match the target conversations. For example, marked by the red cross, PlatoLM has 60% of sample ADV below 15%, while USP has only 5% or less ADV. It demonstrates its stronger capability in preserving the semantic characteristics of the original dialogues. The uniformly lower curves of USP and USP w/o RLCC compared to other baselines (PlatoLM, ProfileGPT(4o), and DialogueGPT(4o)) suggest that our approach generates dialogues that are more faithful to the target conversations across typical and extreme cases.
We also analyze and demonstrate that our user simulator can sample different representatives (majority and minority) of users compared to random sampling in Appendix C.1 and apply it to downstream applications in Appendix C.3.
Figure 3: Cumulative distribution of ADV performance comparison across different models.
6 Analysis
6.1 Ablation Study
To evaluate the relative importance of RLCC’s two rewards, we tested different values of $\lambda$ in Equation 6, denoted as USP($\lambda$: $1-\lambda$).
As shown in Table 6, $\lambda$ = 0.8 provides the optimal balance between maintaining model capabilities and enhancing dialogue consistency. Higher values of $\lambda$ (0.9) compromise speaking style authenticity without improving r-DPC, leading to superficial profile matching, as shown by increased P.Cover scores. Conversely, $\lambda$ = 0.5 achieves authentic style features but lacks sufficient consistency emphasis, resulting in stagnant performance across capabilities.
6.2 Applications: Dynamic Multi-turn Evaluation For LLMs
One application of our simulator is to fill the gap in the current dynamic multi-turn evaluation of large models. The user simulator can simulate different user groups to dynamically interact with the tested model in multiple rounds and reveal their specific defects, as shown in Table 7.
Model Setup Sampling Strategy Avg. Ranking in LiveBench/
Major Minor Virtual Chatbot-Arena
Deepseek-v3 8.25 6.13 7.70 7.36 1
GPT-4o 7.86 6.65 7.19 7.23 3
Claude-Sonnet 7.18 6.61 7.48 7.09 2
4o-Mini 6.84 5.70 5.52 6.02 4
Claude-Haiku 4.88 5.42 5.43 5.24 5
Table 7: Response model performance comparison over different target groups.
We simulated 300 diverse user profiles by USP, comprising 100 highest-probability profiles(the majority), 100 lowest-probability profiles(the minority) based on estimated density, and 100 random synthetic profiles(the virtuality), using the sampler mentioned in Section 4.4. Then, we leverage USP using these profiles to chat with LLM in multi-turns and evaluate them, followed by MT-Bench (Zheng et al., 2024). The experimental results indicate that our user simulator is effective and consistent with the average rankings on the latest Livebench White et al. (2024) and Chatbot-Arena Chiang et al. (2024). In addition, it can also be seen that GPT-4o is better at catering to minority groups than other models, indicating its superior robustness. A more detailed analysis can be found in the appendix C.3.
7 Conclusion
In this work, we propose a novel user simulator with implicit profiles that excels in authenticity, consistency, and diversity. Based on this, we introduce the USP framework, which integrates extracted user profiles into the user simulator by conditional fine-tuning and reinforcement learning with cycle consistency. Our experimental results, validated by both automatic metrics and human evaluations, show that USP significantly outperforms role-playing simulators (e.g., GPT-4o) and direct simulation approaches (e.g., PlatoLM) in authenticity while achieving comparable consistency at both the sentence and conversation levels. Furthermore, through a dynamic evaluation across various LLMs chatting with diverse demographic groups, we demonstrate USP’s effectiveness in real-world applications.
Limitations
We acknowledge the following limitations: 1) Applicability Across Different Scenarios: We conduct the experiments on a single dataset, and there has been limited validation across multiple datasets to assess the generalizability of the results. 2) Cultural and Linguistic Scope: We focus on English dialogues in this paper, which may limit the applicability of USP to other linguistic and cultural contexts.
Ethics Statement
Although LMSYS-1M has undergone extensive data cleaning and ethical checks, the dataset may still contain sensitive or harmful content, reflecting violent, explicit, or discriminatory traits in certain dialogue. This could result in USP generating unsafe dialogues. We strongly advise against including sensitive terms in profiles when using USP, as this may lead to extreme behavior in both the USP and the response model.
Appendix A Dataset Construction
A.1 Preprocessing
Our dataset preprocessing follows the method outlined in PlatoLM Kong et al. (2024), which includes the removal of non-English content, filtering of toxic data, elimination of exact duplicates at the dialogue level, and segmentation of conversations into maximum-length token sequences. To maintain discourse integrity, truncated dialogues are ensured to start with the assistant’s turn, preserving context consistency and coherence.
A.2 Profile Dataset
As discussed in Section 4.1, we categorize all attributes into three types: scene-consistent attributes, scene-related attributes, and deep intrinsic characteristics. For scene-consistent attributes, we use the prompt shown in Figure 13, with each metric definition following the guidelines outlined in Cheng et al. (2024b). For scene-related attributes, we use the prompt in Figure 14, and for deep intrinsic characteristics, we refer to Figure 12. The definition of the Big Five Traits scores follows Gosling et al. (2003).
Next, we concatenate the attributes, remove invalid values (e.g., null or meaningless values), and shuffle the order to eliminate any positional bias in the generated profiles. The attributes, which encompass the three aforementioned aspects, are then rephrased using GPT-4o, with the prompt shown in Figure 15. As a result, we obtain automatically labeled profiles for each data entry. The length statistics are shown in Table 8.
Dataset Train Val Test Profile
LMSYS-USP 1,149 1,295 1,438 231
Table 8: Average token length of LMSYS-USP dataset.
Furthermore, we calculated the frequency of occurrence for each attribute value (i.e., the average number of different attribute values per sample) to assess the prevalence of each attribute. The statistics for the objective facts can be found in Figure 4, while for the subjective characteristics, we focused on whether the Big Five Traits were significantly exhibited. Specifically, we only consider traits with high or low scores, while moderate scores are viewed as the average representation of human behavior Moon et al. (2024) and are not included in the subsequent profiles.
Attribute High Rate (%) Low Rate (%)
Conscientiousness 78.07 7.53
Agreeableness 6.45 14.98
Extraversion 4.08 14.15
Openness 58.77 5.30
Neuroticism 2.04 10.12
Table 9: Summary of extracted subjective attribute statistics.
Figure 4: Frequency of occurrence of values across different attributes of objective facts in the attribute extraction process.
A.3 Resource Consumption in Implementation
During the data construction process, each attribute extraction requires approximately \$0.003 using the GPT API. Since each sample requires three extractions, the cost per sample is approximately \$0.01. With a total of around 94,000 samples, the cost for attribute extraction amounts to approximately \$940. Additionally, rewriting the attributes into profiles incurs a cost of around \$0.05 per sample. Therefore, the total cost for constructing the dataset is approximately \$1,400.
Appendix B Implement Detail
B.1 Trainable Model Setup
We use the llama3-base model as the base architecture. Directly training an instruct model with user simulation contradicts the inherent task requirements during its SFT phase, making convergence difficult. We first perform conditional SFT on the training dataset, followed by our diverse profile sampler, which randomly selects 1,000 samples from the training set for virtual user sampling. Specifically, we combine objective facts and subjective descriptions from different profiles and generate approximately 1 million profiles. From these, we select the 5,000 profiles with the lowest similarity to the training dataset and use them for RLCC phrase. The conditional SFT is conducted using four A100 40GB GPUs for full fine-tuning with epoch set to 3, taking about two days. The RLCC phase is trained using two H20 96GB GPUs over the course of five days.
For the PlatoLM, we also use llama3-base as the base architecture. The system prompt used is: “A chat between a curious human and an artificial intelligence assistant. The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.” We fine-tune the model using four A100 40GB GPUs with epoch set to 3, which takes approximately two days.
For the AI detection model, we follow Cheng et al. and use Longformer Beltagy et al. (2020) to train on our dataset. Since our dataset naturally distinguishes between AI and human-generated text, we label user utterances as human and assistant utterances as AI. We trained for 3 epochs using dual 3090 GPUs, taking three days to complete
For the profile generator model, we utilized LLaMA3-Instruct AI@Meta (2024) as the backbone and fine-tuned it on our carefully curated profile dataset. This approach effectively distilled the two-stage profile generation capabilities of GPT-4o. Training was conducted on four A100 40GB GPUs for 3 epochs, taking two days to complete.
B.2 Baseline Model Setup
Table 10 presents the experimental setup for the baseline models. For models based on GPT, we use the corresponding APIs with default settings for inference. For other models, the experiments are conducted on a single NVIDIA RTX A100 GPU with a batch size of 2 and a repetition penalty of 1.0. An exception is made for CharacterGLM, where we utilize its chat function with a repetition penalty of 1.6 and set the number of beams to 3.
Model Name Backbone System Prompt
USP LLAMA3-8B
You are engaging in a conversation with an AI assistant. your profile is:
{profile}.
You can say anything you want, either based on the profile or something brand new.
DialogueGPT(4o) gpt-4o-2024-08-06 -
PlatoLM LLAMA3-8B
A chat between a curious human and an artificial intelligence assistant. The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.
ProfileGPT(4o) gpt-4o-2024-08-06
You are engaging in a conversation with an AI assistant. your profile is:
{profile}.
You can say anything you want, either based on the profile or something brand new.
CharacterGLM ChatGLM-6B
以下是一段User和AI assistant之间的对话。
关于User的信息:{profile}
关于AI assistant的信息:GPT-4o
Character-LLM-Socrates-7b LLAMA-7B
I want you to act like the person described in the profile below: {profile}. I want you to respond and answer like the person, using the tone, manner, and vocabulary the person would use. You must know all of the knowledge of the person.
The status of you is as follows:
Status: You are engaging in a conversation with an AI assistant.
The interactions are as follows:
Table 10: Baseline model setup.
B.3 Seed conversation design
As there are two types of simulators, one being a response model based on role-playing, which cannot proactively initiate conversations, we first embed the corresponding profile into the system prompt. Then, by using the query "What will you say to start the conversation?" we guide the model to simulate the user’s input over the test dataset. Finally, the simulated user utterance is passed to the response model for interactive generation. The second type of simulator can directly generate reasonable user utterances without requiring additional guidance after setting up the system prompt. Specific configurations of the system prompts for different baseline models are outlined in Table  10.
B.4 Human Evaluation
B.4.1 Profile Evaluation
We employed two annotators to rate the extracted profiles on a scale of 1 to 5 based on the given dialogues, assessing the accuracy and completeness of objective facts, reasonableness of subjective descriptions, and overall naturalness. For accuracy, we emphasized precision without hallucinations; for completeness, we focused on the absence of significant omissions; and for reasonableness, we prioritized rational, well-justified descriptions, avoiding exaggeration or bias. The results in Table 11 show that annotators achieved moderate to high agreement on profile assessment. The evaluation interface is illustrated in Figure 6.
Dataset Profile Source Objective Facts Subjective Characters Naturalness
LMSYS-USP GPT-4o 4.64 4.19 4.66
Table 11: Human evaluation results for the quality of constructed profiles.
B.4.2 Dialogue Evaluation
We recruited eight annotators to assess conversation-level results based on two key criteria: authenticity and consistency. For authenticity, annotators were tasked with identifying user utterances within a dialogue that most closely resemble those in a reference dialogue, evaluating them across three dimensions: Style, Semantics, and Quality. For consistency, annotators were provided only with the user profile and asked to select the user utterances that best align with the profile, considering Accuracy, Completeness, and Quality.The definitions follow prior work Cheng et al. (2024a, b), with detailed guidelines illustrated in Figure 16 and Figure 17.
To minimize position bias and the potential influence of prior exposure, each annotator was randomly assigned pairs of dialogues, with the order of options also randomized. The evaluation interfaces for authenticity and consistency assessments are presented in Figure 7 and Figure 8, respectively.
Figure 5: Distribution of different sampling strategies.
Figure 6: Human evaluation interface for extracted profile evaluation.
Figure 7: Human evaluation interface for assessing dialogue authenticity. Evaluators compare two generated conversations and select which user’s utterances more closely align with those in the reference dialogue.
Figure 8: Human evaluation interface for assessing dialogue consistency. Evaluators determine which user from two test conversations better aligns with a provided user profile.
Appendix C Further Analysis
C.1 Sampling strategy effectiveness
To evaluate our density sampler’s effectiveness, we use two complementary metrics: Local Density Loss (LDL) Rodriguez and Laio (2014) for structure preservation and Uniformity Loss Wang and Isola (2020) for global coverage. Lower LDL values indicate tighter local clustering, confirming preservation of natural profile distribution structures, while lower Uniformity Loss values demonstrate better global coverage while maintaining realistic distributions. Guided by GKDE density distribution, we implement two targeted strategies: sampling high-density regions for majority patterns and weighting low-density regions for minority case coverage. Figure 5 demonstrates how this approach maintains distribution balance while achieving these objectives. Moving right to left on the uniformity loss axis reveals a clear transition from majority to minority sampling patterns. Majority samples (blue circles) perform optimally in low LDL regions with high uniformity, preserving key distribution characteristics. Minority samples (orange squares) strategically occupy higher density regions with moderate uniformity, capturing diverse distribution patterns. The steady progression of sampling percentages in both cases demonstrates controlled sampling behavior, while random sampling (green triangles) shows more scattered patterns, validating our approach’s reliability. The overall performance marker (red star) at the intersection confirms our method successfully balances distribution preservation and targeted sampling objectives.
C.2 Case Study
To evaluate USP’s specific performance in terms of consistency and authenticity, we present two non-cherry-picked dialogues generated through interactive conversations with GPT-4o. Figure 10 examines authenticity by comparing against reference dialogues and other baselines, while Figure 11 evaluates profile consistency among profile-based baseline models.
In terms of authenticity, USP successfully captures both stylistic nuances (such as the consistent use of lowercase "i" and concise questioning patterns) and semantic alignment with target conversations. In contrast, PlatoLM deviates from the dialogue flow by the fourth turn, while ProfileGPT(4o) falls into repetitive appreciation patterns.
Regarding consistency, USP effectively synthesizes both objective profile information and subjective characteristics, demonstrating robust generalization capabilities when handling previously unseen user profiles.
C.3 Downstream analysis
(a) Performance on the majority group
(b) Performance on the minority group
(c) Performance on the virtual group
Figure 9: Performance trends of different LLM models across dialogue turns for various demographic groups.
Our analysis of performance trends across dialogue turns for mainstream LLMs with different demographic groups reveals four key findings, as illustrated in Figure 9: (1) While LLMs demonstrate robust performance with majority demographics, they show notably decreased overall effectiveness when interacting with minority groups, highlighting limitations in personalization capabilities; (2) The models maintain reasonable performance with virtual groups, suggesting effective generalization abilities beyond real-world demographics; (3) Instruction-following capability gradually declines as dialogue turns increase, aligning with observations from previous studies Kwan et al. (2024); Maharana et al. (2024); (4) The pronounced performance volatility across dialogue turns for minority groups underscores the need for enhanced capabilities in processing and responding to less common interaction patterns.
Appendix D Prompt Templates
Case Study on Authenticity of Simulated Dialogue by Different User Simulators ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): Hey there! I’m really excited to dive into some AI discussions with you… Assistant: Hello! I’m thrilled to chat about AI with you… ProfileGPT(4o): It’s great to dive into these with someone who shares such a passion… Assistant: I’m glad to hear that you’re as passionate about AI as I am! Your interests… ProfileGPT(4o): I’d be delighted to guide you in translating these advanced AI concepts… Assistant: I’m excited to help guide your ideas into actionable solutions … ProfileGPT(4o): I’m thrilled at the prospect of collaborating to develop solutions… Assistant: I’m excited at the prospect of working together to make these AI strategies… PlatoLM Simulated Dialogue: PlatoLM: how to train a model like you Assistant: I am a large language model… preprocessing step involves tokenizing the text… PlatoLM: what do you mean by tasks such as tokenizing text… explain these terms please? Assistant: Certainly! These terms refer to common preprocessing steps… PlatoLM: so what does the final product look like when we finish tokenizing? Assistant: Absolutely! Let’s go through a brief example step by step… PlatoLM: who painted NAME_27? Context knowledge: #title NAME_27 … Assistant: Based on the provided context, the artist referred to as… USP Simulated Dialogue: USP: hello who am i talking to? Assistant: Hello! You’re speaking with ChatGPT… USP: are you vicuna? Assistant: No, I’m not Vicuna. I’m ChatGPT, created by OpenAI… USP: what do you know about it? Assistant: Vicuna is an AI language model developed by a different research group… USP: can you tell me how i could train and deploy my own ai model? Assistant: Training and deploying your own AI model involves several key steps… Reference Dialogue: User: how to train a model like you Assistant: I am a large language model, which means that I was trained on… User: what is your base model Assistant: I am a large language model called GPT… User: isn’t you vicuna Assistant: I am not aware of any information or research about a model called "vicuna"… User: what hardware do i need to run you Assistant: I am a large language model called GPT, which was trained on…
Figure 10: Case study comparing USP with other user simulators in the first four turns (out of 10 turns). USP and ProfileGPT(4o) depend exclusively on the provided profile, while PlatoLM utilizes the first-turn context. All simulators interact with GPT-4o for dialogue generation, and the reference dialogue below serves as the target for reconstruction.
Case Study on Consistency of Simulated Dialogues across Different User Simulators Profile: You are engaging in a conversation with an AI assistant. your profile is: You have a passion for engaging in activities that challenge your vocabulary and word formation skills. One of your favorite pastimes is diving into puzzle games that offer a thrilling test of language prowess. You often find yourself thinking of five-letter words with tricky placements, such as words containing ’A’ in the fourth position, or concocting ones with ’R’ in the second spot and ’T’ at the end. This aligns with your appreciation for linguistic puzzles and an enduring enjoyment of wordplay, reflecting your cognitive curiosity and dedication to enhancing your language skills. Your personality shines through a playful and engaging language style, particularly evident when you use structured questions to create dynamic game scenarios. You approach life with a vivid imagination, delighting in intellectual challenges that require intricate problem-solving. Your attention to detail is noteworthy, and you thrive in situations where organization is key. Preferring a more reserved lifestyle, you cherish moments of solitude and introspection, embracing activities that allow you to delve into complex, solitary intellectual pursuits with enthusiasm and precision. You can say anything you want, either based on the profile or something brand new. ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): How about a playful journey into the world of word puzzles? Let’s start with… Assistant: Sure! How about the word "brave"? The letter ’A’ is in the fourth position. ProfileGPT(4o): Wonderful! "Brave" is both intriguing and inspiring. Now, let’s spice it up… Assistant: Absolutely! The word "crest" fits your criteria, with ’R’… ProfileGPT(4o): Splendid! "Crest" evokes such a vivid image; I love it…. Assistant: Of course! The word "shine" fits your criteria, with ’S’ as the first letter … ProfileGPT(4o): Brilliant choice! "Shine" has such a radiant quality to it… Assistant: I’m glad you’re enjoying it! How about this: I’ll give you a word puzzle… USP Simulated Dialogue: USP: Let’s play a game about 5 letter words. I’ll ask you some questions and try to guess the answer. Can we start? Assistant: Sure, I’d be happy to play a game with you! Let’s get started! USP: Question: What is a 5-letter word which has "A" in its 4th place? Answer: Assistant: Is the word you are thinking of "Polar"? USP: Nope What about "gnaw"? Assistant: "GNW" is a four-letter word. Let’s think of another five-letter word with "A" in the… USP: Next Question: Give me one example of a 6-letter word having R at second place and T at sixth place. Assistant: Sure! A six-letter word with "R" in the second position and "T" at the sixth position… USP: Great! Now give me one example of a 7-letter word having S at third place and H at seventh place. Assistant: Sure! One example of a seven-letter word with "S" in the third place and "H" in …" Reference Dialogue: User: You are a very intelligent puzzle solver. We are going to play a game…Question: What is a 5 letter word which has "A" in 4th position? Answer: Assistant: Panda. User: Question: Suggest a five letter word which has R in second position and T in fifth position. Answer: Assistant: Graft.
Figure 11: Case study of user simulator comparisons showing the first four turns of a 10-turn dialogue. USP and other simulators interact with GPT-4o using only the provided profile, with the reference dialogue serving as the reconstruction target.
Prompt for Extracting Big Five Personality Traits Task: Analyze the provided dialogue to assess the user’s personality traits across 5 personality dimensions. Focus exclusively on the user’s characteristics, disregarding any information related to others individuals, unless it directly impacts the user. For each personality trait: 1. Assign each dimension independently with score: "High", "Low," or "Inconclusive." 2. Provide conclusion: A high-level description, with concise supporting details. 3. Provide reason: Justify the assigned score with specific evidence from the dialogue. 4. Mark traits as Inconclusive only when no clear evidence exists. Personality Trait Defination: {{metric}}: {{definition}} Format:
    "Trait Name": {
        "score": "High/Low/Inconclusive",
        "conclusion": "The user is a [trait descriptor] person...",
        "reason": "Explanation referencing specific dialogue evidence."
Example:
[User]: "She is my age, in a homeless women’s shelter, living under very poor
conditions. She is a mental health client, but the treatment team seems to
... Her background is similar to mine, and I cannot abandon her."
Detected Personality Traits:
  "Conscientiousness": {
    "score": "High",
    "conclusion": "The user is a conscientious person who demonstrates a
    sense of duty and commitment.",
    "reason": "The user expresses a strong sense of responsibility ..."
  "Agreeableness": {
    "score": "High",
    "conclusion": "The user is an empathetic and compassionate person who
    values relationships.",
    "reason": "The user shows care and concern for their cousin’s well-being..."
  "Extraversion": {
    "score": "Inconclusive"
Figure 12: Prompt for extracting deep intrinsic characteristics.
Prompt for Extracting Scene-Consistent Attributes User Persona Analysis Task Objective The primary goal of this task is to analyze user utterances in-depth and accurately extract key persona attributes based on both direct and implicit cues. These attributes should be categorized into distinct fields, with any missing or unclear details left blank. Field Descriptions: {{Field}}: {{definition}} Guidelines 1. Carefully examine each user utterance to extract relevant persona traits. Consider both direct statements and implicit clues. 2. Ensure that the extracted attributes are specific and directly relevant to the user’s utterances. Avoid vague or generalized descriptions unless explicitly supported by the text. 3. Pay attention to distinctive communication styles (e.g., formal or casual tone, frequent use of specific words or phrases) to capture the user’s unique way of communicating. Example User Utterances:
[User]: Given an array of integers nums and an integer target, return indices of
the two numbers such that they add up to target in Python...
[Assitant]: To solve the problem...
[User]: Thanks
Expected Output:
    "gender": [],
    "age": [],
    "location": [],
    "occupation": [
        "Likely a beginner programmer or student studying computer science,
        as evidenced by the simple coding problem in Python."
    "education": [
        "Possibly a student in computer science or a related field,
        at an early stage in learning programming, specifically Python."
    "family_relationships": [],
    "routines_or_habits": [],
    "social_relationships": [],
    "language_style": [
        "Concise and task-oriented",
        "Polite response ’Thanks’ after getting satisfactory answer"
    "other_experiences": []
Figure 13: Prompt for extracting scene-consistent attributes.
Prompt for Extracting Scene-Related Attributes User Persona Analysis Task Objective The goal of this task is to analyze multi-turn user utterances within a conversation with an assistant and extract key elements such as the primary goals and specific task descriptions. Each extracted detail should be as specific as possible, reflecting the user’s context, objectives, and scenario. Output Format The extraction should be presented in a structured JSON format, as shown below:
    "scenarios": [
            "goals_or_plans": "<List of User’s goals or plans>",
            "task_details": "<List of specific tasks summary made by the user>"
Field Descriptions: goals_or_plans: User’s short-term or long-term objectives, derived from explicit statements or inferred from the overall conversation. If no explicit goals are stated, infer them from the main topics discussed. task_details: Specific tasks, actions, or requests made by the user. Each task should be concisely summarized with specifics. If there are multiple tasks or requests, list each one separately. Example User Utterances:
[User]: Summarize: Harry Potter is a fictional character in the Harry Potter series...
[Assitant]: Harry Potter is a fictional character...
[User]: Write an email inquiring about coursework...
    "scenarios": [
            "goals_or_plans": "Aiming to gain a deeper understanding of the Harry
                        Potter series, possibly for academic or personal enrichment.",
            "task_details": [
                "Summarizing introductory content about the Harry Potter character."
            "goals_or_plans": "Looking to improve professional communication skills.",
            "task_details": [
                "Writing an email to inquire about coursework."
Figure 14: Prompt for extracting scene-related attributes.
Prompt for Rephrasing Attributes into Natural Descriptions Narrative Generation Objective Rephrase the provided key-value pairs into a natural, coherent second-person description. Core Requirements 1. Perspective: Use second-person perspective ("you", "your"). 2. Structure: Two paragraphs: First paragraph: Present objective facts. Second paragraph: Describe subjective characteristics. 3. Key Principles Accurately represent all provided details. Ensure the language flows naturally, remains engaging, and avoids redundancy. Focus on clear and seamless transitions between ideas. Output Expectations Objective Facts: Convert the key-value pairs into a clear and natural description without over-explaining or adding unnecessary details. Ensure each scenario is logically connected and key information is presented effectively. Subjective Characteristics: Avoid vague terms like "high perfectionism" or "moderate emotional stability." Use vivid, descriptive language to bring these traits to life.
Figure 15: Prompt for rephrasing attributes into natural descriptions for profile generation.
Human Evaluation Guidelines for Authenticity 1. Task Description: Please choose which user in the two test conversations is more similar to the reference conversation being spoken by the same person. 2. Evaluation Criteria: Semantic Similarity: Measure the thematic consistency and discourse coherence between the generated user utterance and the target user utterance. Preference should be given to the utterance that more accurately reflects the semantic content of the target. Stylistic Parity: Analyze whether the generated user utterance matches the style of the target user utterance, including its tone, vocabulary, and grammatical structure. The utterance that aligns more closely with the stylistic elements of the target should be favored. Quality: Examine the fluency and logical coherence of the user utterance, focusing on the linguistic and logical smoothness of the user utterance. The more coherent and fluent utterances should be chosen.
Figure 16: Human evaluation guidelines for authenticity.
Human Evaluation Guidelines for Consistency 1. Task Description: Select the dialogue that contains the most appropriate user utterance from the two given generated dialogues based on the target user profile. 2. Evaluation Criteria: Persona Reflection: Evaluate how closely each user utterance reflects the target profile’s thematic content, tone, and stylistic features. Preference should be given to the dialogue whose user utterance most accurately reflects the profile’s characteristics in terms of thematic coherence and expression. Comprehensiveness: Assess the extent to which the user utterance encapsulates the target profile, integrating both objective facts and subjective descriptions. The more comprehensive utterance, which addresses a broader range of profile aspects, should be favored over one that focuses narrowly on a single dimension. Quality: Evaluate the fluency, coherence, and human-likeness of the user utterance. Preference should be given to the utterance that demonstrates greater linguistic smoothness, logical cohesion, and alignment with genuine human conversational patterns.
Figure 17: Human evaluation guidelines for consistency.
Prompt for NLI Score of Profile Precision Based on Given Dialogue Role You are an expert in evaluating the consistency between a given user profile (Source) and the user’s utterance (Target). Your task is to assess whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
  "score": <score>,
  "reason": "<concise explanation of the reasoning>"
Example: Source: You are interested in dataset-related details. Target: [User]: Show me how to implement a toy version of a relational database. Output:
  "score": 1,
  "reason": "The request for implementing a relational database suggests an
  interest in data structures and datasets, which aligns with the Source."
Guidelines: 1. Contextual Inference: Do not evaluate the Target in isolation. Instead, determine its logical relationship to the Source, considering both explicit statements and reasonable implications. 2. Concise & Precise Justification: The reasoning should be clear, objective, and free from unnecessary elaboration.
Figure 18: Prompt for DP.P based on NLI.
Prompt for NLI Score of Dialogue Precision Based on Given Profile Role You are an expert in evaluating consistency between a given dialogue history (Source) and a corresponding user profile description (Target). Your task is to determine whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
  "score": <score>,
  "reason": "<concise explanation of the reasoning>"
Example : Source: User (Turn-1): Show me how to implement a toy version of a relational database. User (Turn-2): Thanks a lot! Target: You are polite. Output:
  "score": 1,
  "reason": "The user’s expression of gratitude in Turn-2 indicates politeness,
  which aligns with the Target."
Guidelines: 1. Contextual Inference: Do not evaluate the Target in isolation. Instead, determine its logical relationship to the Source, considering both explicit statements and reasonable implications. 2. Concise & Precise Justification: The reasoning should be clear, objective, and free from unnecessary elaboration.
Figure 19: Prompt for DP.R based on NLI.
Prompt for Subjective Characteristics Score (SC.Score) in Consistency Evaluation Task Description You are tasked with evaluating the quality of user responses in real human-LLM interactions. Specifically, you will assess the degree to which a given response (Target) aligns with a predefined personality profile, tone, and linguistic characteristics (Source). Additionally, you must consider the naturalness and authenticity of the Target, ensuring it reflects genuine human conversational patterns. Evaluation Criteria Your assessment will focus on two primary dimensions: 1. Human-Likeness: The extent to which Target exhibits natural human language, characterized by appropriate syntax, tone, and conversational flow. 2. Alignment with Source: The degree to which the Target adheres to the personality traits, tone, and linguistic features specified in the Source. Scoring Guidelines Assign a score from 1 to 5 based on the following criteria: Score 5: The Target fully aligns with the Source and demonstrates exceptional human-likeness. Score 3: The relationship between the Target and Source is ambiguous or unclear, lacking sufficient evidence for alignment or contradiction. Score 1: The Target significantly deviates from the Source or lacks human-likeness, rendering it unnatural or inconsistent. Output Requirements Provide your evaluation in the following JSON format:
  "score": <score>,
  "reason": "<concise reason>"
Key Considerations 1. Contextual Inference: Analyze both explicit content and implicit nuances in the Target to determine its alignment with the Source. 2. Conciseness and Precision: Ensure that your reasoning is clear, objective, and free of superfluous elaboration. 3. Human-Likeness Emphasis: A lack of human-likeness, even if the Target aligns with the Source, will result in a lower score.
Figure 20: Prompt for evaluating consistency in subjective characteristics.
Prompt for Validation Score (Val.Score) in Assessing the Quality of Subjective Characteristics in Profiles Role As an expert in evaluating the consistency between user utterances in a dialogue (Source) and a provided subjective description (Target), your task is to assess whether the Target accurately reflects the characteristics described in the Source, including personality traits, tone, and other relevant attributes. You will then rate this consistency on a scale from 1 to 5. Task Instructions For each pair of Source-Target, apply the following scoring criteria to determine their relationship: Score 5: The Target completely aligns with the Source, with no discrepancies. The profile perfectly represents the characteristics observed in the user’s utterance. Score 3: Ambiguity or insufficient evidence exists, making it difficult to ascertain the relationship definitively. Score 1: A clear discrepancy exists, with the Target significantly deviating from the Source, indicating a mismatch in the represented characteristics. Output Format Your assessment should adhere to the following structured JSON format:
  "score": "<numerical score>",
  "reason": "<a succinct explanation providing justification for assigned score>"
Guidelines: 1. Contextual Inference: Determine the target’s logical relationship to the Source, considering both explicit statements and reasonable implications. 2. Concise & Precise Justification: The reasoning should be clear, objective, and free from unnecessary elaboration.
Figure 21: Prompt for validation score (Val.Score) in assessing the quality of subjective characteristics in profiles.

### v3

@@
 User simulators are crucial for replicating human interactions with dialogue systems, supporting both collaborative training and automatic evaluation, especially for large language models (LLMs).
-However, existing simulators often rely solely on text utterances, missing implicit user traits such as personality, speaking style, and goals.
-In contrast, persona-based methods lack generalizability, as they depend on predefined profiles of famous individuals or archetypes.
-To address these challenges, we propose User Simulator with implicit Profiles (USP), a framework that infers implicit user profiles from human-machine conversations and uses them to generate more personalized and realistic dialogues.
-We first develop an LLM-driven extractor with a comprehensive profile schema.
-Then, we refine the simulation through conditional supervised fine-tuning and reinforcement learning with cycle consistency, optimizing it at both the utterance and conversation levels.
-Finally, we adopt a diverse profile sampler to capture the distribution of real-world user profiles.
-Experimental results demonstrate that USP outperforms strong baselines in terms of authenticity and diversity while achieving comparable performance in consistency.
-Furthermore, dynamic multi-turn evaluations based on USP strongly align with mainstream benchmarks, demonstrating its effectiveness in real-world applications  [^1] All resources will be released at https://github.com/wangkevin02/USP.
+However, current role-playing methods face challenges such as a lack of utterance-level authenticity and user-level diversity, often hindered by role confusion and dependence on predefined profiles of well-known figures.
+In contrast, direct simulation focuses solely on text, neglecting implicit user traits like personality and conversation-level consistency.
+To address these issues, we introduce the User Simulator with Implicit Profiles (USP), a framework that infers implicit user profiles from human-machine interactions to simulate personalized and realistic dialogues.
+We first develop an LLM-driven extractor with a comprehensive profile schema, then refine the simulation using conditional supervised fine-tuning and reinforcement learning with cycle consistency, optimizing at both the utterance and conversation levels.
+Finally, a diverse profile sampler captures the distribution of real-world user profiles.
+Experimental results show that USP outperforms strong baselines in terms of authenticity and diversity while maintaining comparable consistency.
+Additionally, using USP to evaluate LLM on dynamic multi-turn aligns well with mainstream benchmarks, demonstrating its effectiveness in real-world applications.
+We open-source related resources in https://github.com/wangkevin02/USP.
 {CJK}
@@
 Modeling Human-Like User Simulators via Implicit Profiles
-Kuang Wang1,2, Xianfei Li1, Shenghao Yang1, Li Zhou1, Feng Jiang1,2,3 thanks: Feng Jiang is the corresponding author., Haizhou Li1,2 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 3Shenzhen University of Advanced Technology kuangwang@link.cuhk.edu.cn, jeffreyjiang@cuhk.edu.cn
+Kuang Wang1, Xianfei Li1, Shenghao Yang1, Li Zhou1, Feng Jiang2,1thanks: Feng Jiang is the corresponding author and Shenzhen University of Advanced Technology is the corresponding affiliation., Haizhou Li1,3 1SRIBD, School of Data Science, The Chinese University of Hong Kong, Shenzhen, Guangdong 2Artificial Intelligence Research Institute, Shenzhen University of Advanced Technology 3Department of ECE, National University of Singapore kuangwang@link.cuhk.edu.cn, jiangfeng@suat-sz.edu.cn
 1 Introduction
-Figure 1: Examples of performance across various user simulators in multi-turn human-LLM interactions. OF and SC refer to objective facts and subjective characteristics, respectively.
+Figure 1: Examples of different user simulators in multi-turn human-LLM interactions.
+"OF" and "SC" represent objective facts and subjective characteristics, respectively.
+highlights inconsistencies with the target profile, while indicates inauthentic user imitation.
 The user simulator is designed as a proxy for real users in interactions with large language models (LLMs).
-It can simulate realistic user by generating the target user’s behavior or utterances based on the specified characteristics, enabling dynamic multi-turn interactions with LLMs Wan et al.
+It can simulate a realistic user by generating the target user’s behavior or utterances based on the specified characteristics, enabling dynamic multi-turn interactions with LLMs Wan et al.
 (2022) and scene reproduction Wang et al.
@@
 (2024).
-Recent LLMs advance promote user simulators by enhancing naturalness and utility Deng et al.
+Recent advances in LLMs have spurred the development of user simulators, improving their naturalness and utility Deng et al.
 (2024); Zhang et al.
+(2024a).
+Mainstream LLM-based role-playing methods Moon et al.
+(2024) use predefined profiles to mimic diverse user traits.
+However, as LLMs are typically trained to be universally polite and helpful Lu et al.
+(2024), they often lack utterance-level authenticity and struggle with role confusion between user simulation and their inherent assistant nature Xu et al.
 (2023a), as shown in Figure 1.
-Since directly using LLM as the user simulator suffers from role confusion Xu et al.
-(2023a), some works Xu et al.
-(2023a); Kong et al.
-(2024); Sun et al.
-(2024) attempt to enhance its authenticity by training a user simulator on the conversation dataset.
-However, they are only trained on text utterance, making it difficult to simulate diverse user behaviors without seed context, exhibiting limited self-awareness Tseng et al.
-(2024) and failing to maintain a consistent personality.
-Although some LLM-based role-playing methods Moon et al.
-(2024) utilize predefined profiles to alleviate this problem, they require additional extensive annotations and can only be applied to celebrities, lacking the diversity of simulation.
-To address the issues above, we believe that a user simulator knows users’ intrinsic characters hidden in their conversations first and then can be a better simulation.
+Models like PlatoLM Kong et al.
+(2024) and Parrot Sun et al.
+(2024) address this by training on real human-LLM conversation datasets, but they focus only on text utterances, lacking dialogue control.
+This limits conversation-level consistency and diverse simulation without seed context.
+Additionally, they both fail to capture the authentic distribution of user-level diversity, crucial for analyzing group behavior.
+To address the issues above, we believe that a user simulator knows users’ intrinsic characters hidden in their conversations first, and then can provide a better simulation.
 Therefore, we treat user simulation as a dialogue reconstruction task and propose a novel framework named the User Simulator with Implicit Profile (USP).
@@
 (2024b); Tu et al.
-(2024), we then polish the profile attributes into natural, descriptive profiles to ensure generalization.
+(2024) using attributes as profiles, we further polish the profile attributes into natural, descriptive profiles to ensure generalization.
 Then, we integrate the extracted user profiles into the user simulator through two-stage training: (1) conditional supervised fine-tuning with user profiles for utterance-level simulation, and (2) reinforcement learning with cycle consistency to align reflected profiles from simulated dialogues with given profiles for conversation-level simulation.
 We also implement a diverse profile sampler to capture authentic user distributions.
-Our experiments reveal that USP enhances semantic and stylistic similarity of reconstructed multi-turn dialogues to target dialogues by about 34% and 43% over the leading baseline, with reconstruction errors halved, showcasing superior authenticity and diversity.
-It matches the consistency of ProfileGPT(4o), a GPT-4o-based simulation baseline, boosting dialogue profile consistency by 14% in multi-turn scenarios while performing comparably in single-turn cases.
-Furthermore, our USP-based multi-turn dynamic evaluation of LLMs for downstream tasks aligns well with established benchmarks,allowing for a more granular assessment of LLM performance across different user groups.
+Our experiments demonstrate that USP improves semantic and stylistic similarity in reconstructed multi-turn dialogues by approximately 34% and 43% compared to the leading baseline, with reconstruction errors reduced by half, showcasing enhanced authenticity and diversity.
+It achieves dialogue profile consistency comparable to GPT-4o (User w/ Profile), improving multi-turn consistency by 14% while matching single-turn performance.
+Additionally, USP-based multi-turn dynamic evaluation of LLMs for downstream tasks aligns closely with established benchmarks, enabling finer-grained assessment of LLM performance across diverse user groups.
 Our key contributions are outlined below:
   • We propose a novel approach for constructing user simulators using implicit user profiles embedded in human-LLM conversations.
-  • We develop a new framework that infers implicit user profiles as insight, further enhanced with conditional fine-tuning and reinforcement learning with cycle consistency for better simulation at both the utterance and conversation levels.
-  • Experimental results show USP surpasses all baselines in authenticity and diversity, maintains comparable consistency, and proves effective for multi-turn dynamic evaluation of LLMs.
+  • We propose a framework that infers implicit user profiles, further enhanced by conditional fine-tuning and reinforcement learning with cycle consistency, to improve simulation at both the utterance and conversation levels.
+  • Experiments show that USP outperforms baselines in authenticity and diversity, maintains comparable consistency, and enables effective multi-turn dynamic evaluation of LLMs.
 2 Related Works
 2.1 General User Simulator
-Early user simulators including agenda-based methods Schatzmann et al.
+Early user simulators focused on limited action prediction using agenda-based Schatzmann et al.
 (2007); Schatzmann and Young (2009) and model-based methods Asri et al.
 (2016); Kreyssig et al.
+(2018), constrained by early natural language generation capabilities—for instance, generating synthetic binary preferences in conversational recommendation systems Christakopoulou et al.
 (2016).
-These simulators were initially designed with a narrow scope due to limited natural language generation capabilities, such as generating synthetic binary preference responses Christakopoulou et al.
-(2016) in conversational recommendation systems.
 Recent advancements in LLMs enabled more sophisticated simulations of realistic conversations, offering significantly enhanced natural language flexibility.
@@
 2.2 Persona-based User Simulator
-General user simulators often struggle to capture the full spectrum of diverse user needs, leading to a growing interest in persona-based personalization to improve both controllability and diversity in simulations Takanobu et al.
+Since general user simulators often struggle to capture the full spectrum of diverse user needs, it leads a growing interest in persona-based personalization to improve both controllability and diversity in simulations Takanobu et al.
 (2020).
@@
 (2019) to guide user simulators in task-oriented dialogue (ToD) systems.
-With the rise of LLMs and their impressive zero-shot role-playing abilities Njifenjou et al.
-(2024), prompt-driven user simulation has become the dominant approach.
-For example, LLMs have been used with carefully designed predefined profiles to align with human beliefs Chuang et al.
-(2024), simulate consultation scenarios with users exhibiting varying personalities and needs in ToD systems Zhang et al.
-(2024a), and model user preferences in conversational recommendation systems Yoon et al.
+With the rise of LLMs and their strong zero-shot role-playing capabilities Njifenjou et al.
+(2024), prompt-driven user simulation has become the dominant paradigm.
+LLMs have been used to simulate users with predefined profiles Chuang et al.
+(2024), model diverse personalities and needs in ToD Zhang et al.
+(2024a), and capture user preferences in conversational recommendation Yoon et al.
 (2024).
+Our method follows this line of work, with a focus on addressing authenticity, consistency, and diversity, which remain underexplored in related studies.
 3 Task Definition
-We formulate user simulation as a dialogue refactoring task, aiming to replicate multi-turn user behavior in target dialogues.
-Given a target dialogue $d_{i}=\{(u_{i1},r_{i1}),\dots,(u_{ij},r_{ij})\}$ between a user $U_{i}$ and a response model $R_{i}$, where $u_{ij}$ and $r_{ij}$ represent the $j$-th turn user utterance and the corresponding model response, respectively.
-To achieve high-fidelity simulation of user responses within a given context, we aim to minimize the utterance-level distance $D_{\text{utt}}(u_{ij},u_{ij}^{\prime})$ where $u_{ij}^{\prime}\sim P_{prob}(\cdot|c_{ij}^{\prime},U_{i}^{\prime})$.
-Here, $c_{ij}^{\prime}$ represents the context used by the user simulator $U_{i}^{\prime}$ to generate $u_{ij}^{\prime}$.
-This optimization ultimately leads to reducing the dialogue-level distance, as formulated in Eq.
-1.
-$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime})}D_{\text{dia}}(d_{i},d_{i}^{\prime})$ (1)
-where $D_{\text{dia}}$ represents a distance function that evaluates the user utterances between the simulated dialogue $d_{i}^{\prime}$ and the real conversation $d_{i}$.
-Recent studies show that role-playing with specific user profiles ($P_{i}$) can effectively achieve diverse user simulations Liu et al.
-(2023).
-However, unlike celebrities or well-known characters, user profiles in real-world conversations are often implicit and difficult to obtain Wang et al.
+We formulate user simulation as a dialogue refactoring task to replicate multi-turn user behavior in a target dialogue $d_{i}=\{(u_{i,1},a_{i,1}),\dots,(u_{i,n},a_{i,n})\}$, where $u_{i,j}$ is the $j$-th user utterance and $a_{i,j}$ is the corresponding response answer.
+Our goal is to achieve high utterance-level and dialogue-level fidelity.
+Formally, we minimize utterance-level distance $D_{\text{utt}}(u_{i,j},u_{i,j}^{\prime})$ and dialogue-level distance $D_{\text{dia}}(d_{i},d_{i}^{\prime})$, where $u_{i,j}^{\prime}$ is the simulated utterance and $d_{i}^{\prime}$ is the simulated dialogue.
+Direct simulation struggles to capture personalized traits.
+Recent studies Deng et al.
+(2024); Kong et al.
+(2025) demonstrate that role-playing with specific user profiles ($p_{i}$) effectively enables diverse user simulations.
+However, unlike well-known figures, user profiles in real-world conversations are often implicit and challenging to derive Wang et al.
 (2024a).
-To address this, we reformulate the task by first extracting the implicit user profile from the given dialogue using profile extractor $P_{\text{extractor}}$, and then reconstructing a closer dialogue, with an emphasis on the user’s utterances, as described in Eq.
-2.
-$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime},P_{i})}D_{\text{dia}}(d_{i},d_{i}^{\prime}),$ (2)
-where $P_{i}=P_{\text{extractor}}(d_{i})$.
+To address this, we reformulate the task by extracting the implicit user profile from the dialogue using a profile extractor $P_{\text{extractor}}$, then reconstructing the target dialogue as Eq. 1.
+$\min_{\begin{subarray}{c}d_{i}^{\prime}\sim P(\cdot|p_{i},\pi_{\theta})\\ u_{i,j}^{\prime}\sim P(\cdot|c_{i,j},p_{i},\pi_{\theta})\end{subarray}}\left[D_{\text{utt}}(u_{i,j},u_{i,j}^{\prime})+\alpha D_{\text{dia}}(d_{i},d_{i}^{\prime})\right],$ (1)
+where $p_{i}=P_{\text{extractor}}(d_{i})$, $\pi_{\theta}$ represents the learnable parameters of the language model, and $c_{i,j}=\{(u_{i,1},a_{i,1}),\dots,(u_{i,j-1},a_{i,j-1})\}$ denotes the ground-truth context up to the $j$-th turn.
+The hyperparameter $\alpha$ balances the utterance-level and dialogue-level distances.
 4 Modeling User Simulator with Implicit Profiles
 Figure 2: Overview of our proposed User Simulator with implicit Profile(USP) framework.
-To accomplish this task, we propose the User Simulator with Implicit Profiles (USP) framework, as illustrated in Figure 2, which aims to minimize the objective in Eq.
-2 while ensuring authenticity, consistency, and diversity.
+We propose the User Simulator with Implicit Profiles (USP) framework, shown in Figure 2, to minimize the objective in Eq.
+1 across four stages.
+First, we build a user profile extractor with a tailored schema (Section 4.1).
+Then, we optimize utterance-level authenticity using conditional SFT (Section 4.2), ensure conversation-level consistency via Reinforcement Learning with Cycle Consistency (RLCC) (Section 4.4), and achieve user-level diversity through corpus distribution fitting (Section 4.3).
 4.1 User Profile Construction
@@
 (2024b), we design a user profile schema containing objective facts(OF) and subjective characteristics(SC) to represent them, as shown in Table 1.
-The OF focuses on common topics in human conversation (Cheng et al., 2024b; Dunbar et al., 1997) including Scene-Consistent Attributes (such as age, gender, and location) and Scene-Related Attributes (such as goal,and task details).
-SC considers both external and internal personality dimensions represented by language style Wang et al.
-(2024a) and the Big-Five Traits Different from previous work Cheng et al.
+The OF focuses on common topics in human conversation (Cheng et al., 2024b; Dunbar et al., 1997), including scene-consistent attributes (such as age, gender, and location) and scene-related attributes (such as the goal and task details).
+SC encompasses external personality dimensions, reflected in language style Wang et al.
+(2024a), and internal personality dimensions, captured by the Big-Five Traits Gosling et al.
+(2003).
+Unlike prior work Cheng et al.
 (2024b); Tu et al.
-(2024), we reformulate discrete attributes into coherent narrative descriptions to achieve greater generalization and flexibility.
+(2024) that relies on discrete attributes for user profiles, we further reformulate these attributes into coherent narrative descriptions to enhance generalization and flexibility.
 4.1.2 User Profile Extractor
-To obtain such a user profile, we design an LLM-driven user profile extractor extracting the implicit user profile from the human-LLM conversation.
+To obtain such a user profile, we design an LLM-driven user profile extractor that extracts the implicit user profile from the human-LLM conversation.
 The extractor first leverages advanced LLM (such as GPT-4o) to extract the user character attributes mentioned above with a well-designed prompt.
-Then, the extractor collects the valid attributes (No empty) together and polishes them into natural language descriptions.
+Then, the extractor collects the valid attributes together and polishes them into natural language descriptions.
 Further prompt details regarding the extractor can be found in Appendix A.2.
 4.1.3 Profile Quality Verification
-Due to the reliance on predefined user profiles in existing role-playing approaches Zhou et al.
-(2024b), the correlation between user profiles and conversations has been largely overlooked.
-To address this limitation, we propose Dialogue Profile Consistency (DPC) for evaluating OF based on atomic fact verification that measures both precision (DP.P) and recall (DP.RJandaghi et al.
+Existing role-playing methods rely on predefined profiles and dialogues from separate sources, either extracted from novel segments or generated by LLMs Zhou et al.
+(2024b); Chen et al.
+(2023); Shao et al.
+(2023) without verifying alignment Wang et al.
+(2024a); Cheng et al.
+(2024b).
+This overlooks the correlation between profiles and dialogues, potentially hindering simulation performance Yu et al.
 (2024).
-Drawing inspiration from FactScore Min et al.
-(2023) and Con.Score Madotto et al.
-(2019), we first introduce Factual Consistency (Fact.Con), as defined in Eq.
-3.
-Given a target $T$, we evaluate the consistency between the source and target by decomposing $T$ into atomic facts $af_{k}$ using an atomic fact generator (afg).
+To provide an automatic metric for evaluating this, we propose Dialogue Profile Consistency (DPC), which frames consistency as a retrieval task Jandaghi et al.
+(2024).
+DPC employs an F1-score approach, assessing consistency through atomic fact verification by measuring both precision (DP.P) and recall (DP.R).
+We first introduce Factual Consistency (Fact.Con), an adaptation of FactScore Min et al.
+(2023) tailored for dialogue scenarios, as defined in Eq.
+2.
+Given a target $T$, we assess its consistency with the source by decomposing $T$ into atomic facts $af_{k}$ using an atomic fact generator (afg).
 We then compute the natural language inference (NLI) score for each atomic fact with respect to the source $S$.
 […]
+where $\text{NLI}(\cdot,\cdot)$ denotes the NLI model, implemented using prompt-based GPT-4o.
 […]
-Additionally, we use a validation score (Val.Score) to assess the quality of SC descriptions based on the dialogue, rating them on a scale from 1 to 5 using GPT-4o. Detailed prompts in Appendix D.
+Additionally, we use a Validation Score (Val.Score) to evaluate SC description quality based on the dialogue, rated on a 1–5 scale using GPT-4o(prompts detailed in Appendix D).
 4.2 Conditional Supervised Fine-Tuning
 To empower the LLM with the general capability to simulate diverse users at the utterance level, we utilize conditional supervised fine-tuning based on user profiles.
-It enables the LLM to learn the conditional generation mapping based on both the extracted profile $P_{i}$ and context $c_{ij}^{\prime}$.
-As subtle misalignment between the core objectives of the user simulator and the response model, the SFT language modeling loss is adjusted as shown in Eq.
+It enables the LLM to learn the conditional generation mapping based on both the extracted profile $p_{i}$ and context $c_{i,j}$.
+As a subtle misalignment between the core objectives of the user simulator and the response model, the SFT language modeling loss focuses on optimizing user utterance as shown in Eq.
+3.
+$\min_{\pi_{\theta}}\sum_{i,j,k}-\log P(u_{i,j,k}|u_{i,j,<k},c_{i,j},p_{i},\pi_{\theta})$ (3)
+where $u_{i,j,k}$ denotes the $k$-th token of the $u_{i,j}$.
+4.3 Diverse Profile Sampling
+We propose Diverse Profile Sampling to generate naturalistic user profiles that reflect real-world characteristic distributions.
+Our method first embeds constructed profiles into a semantic space using SimCSE Gao et al.
+(2021), followed by dimensionality reduction via UMAP McInnes et al.
+(2018).
+We then apply Gaussian Kernel Density Estimation (GKDE) to fit the underlying distribution, allowing probabilistic sampling of realistic profiles for downstream tasks such as majority representation.
+To further enhance diversity, we synthesize virtual profiles by combining OF and SC descriptions from nearest neighbors, producing novel yet plausible profile variants.
+4.4 Reinforcement Learning with Cycle Consistency
+The conditional SFT stage enables the user simulator to generate human-like utterances focused on forward consistency, producing precise responses aligned with the target profile.
+However, it does not ensure full reflection of the profile, i.e., backward consistency or profile recall.
+To overcome this and improve conversation-level consistency, we introduce Reinforcement Learning with Cycle Consistency (RLCC), which enhances alignment between the user simulator’s actual behavior reflected in simulated dialogues and the target behavior defined by the profile, ensuring a closer match to the intended target profile.
+In this stage, the user simulator $u_{i}^{\prime}$ interacts with a response LLM based on a target profile $p_{i}$, sampled via Diverse Profile Sampling, to generate a simulated dialogue $d_{i}^{\prime}$.
+The dialogue ends when it reaches the maximum context length or a predefined turn limit (set to 10).
+The reflected profile $p_{i}^{\prime}$ is then extracted from $d_{i}^{\prime}$ using the profile generator.
+Our goal is to maximize the semantic similarity between the target profile $p_{i}$ and the reflected profile $p_{i}^{\prime}$, both in objective facts and subjective characteristics, as defined in Eq.
 4.
-$l_{\text{lm}}=\sum_{j}\sum_{k}-\log P_{prob}(u_{i,j,k}^{\prime}|u_{i,j,<k}^{\prime},c_{i,j}^{\prime},P_{i})$ (4)
-where $u_{ijk}$ represents the $k$-th token of the $j$-th utterance from the $i$-th user.
-4.3 Reinforcement Learning with Cycle Consistency
-To further enhance conversation-level consistency, we introduce Reinforcement Learning with Cycle Consistency (RLCC), which optimizes the user simulator by aligning the reflected profile, extracted from simulated dialogues, with the target profile.
-In this stage, we extract the simulator profile $P_{i}^{\prime}$ from dialogues $D^{\prime}$ generated by the user simulator $U_{i}^{\prime}$ using the profile generator, based on virtual profiles sampled by our diverse profile sampler (see Section 4.4).
-Our goal is to maximize the semantic similarity between the target profile $P_{i}$ and the extracted reflected profile both in objective facts and subjective characteristics.
-The dialogue-level reward is then distributed to each user utterance within the dialogue, as shown in Eq.
-5.
-$r^{cc}_{i,j}=\text{sim}(P_{i},P^{\prime}_{i})$ (5)
-Then, we optimize profile recall through Proximal Policy Optimization (PPO) Schulman et al.
-(2017) using cycle consistency as a dialogue-level reward signal.
-It enhances the user simulator’s self-expression in dialogues, moving beyond simple profile adherence or precision consistency.
-Additionally, to prevent reward hacking, we incorporate an AI detection model as an auxiliary reward. The final reward is formulated in Eq. 6.
+$\max_{\pi_{\theta}}\;\mathbb{E}_{p_{i}\sim D,d_{i}^{\prime}\sim\pi_{\theta}(p_{i})}\left[\text{Sim}(p_{i},P_{extractor}(d_{i}^{\prime}))\right]$ (4)
+where $\text{Sim}(\cdot,\cdot)$ is a similarity model (SimCSE Gao et al.
+(2021)), and $D$ denotes the virtual profiles dataset from the sampler in Section 4.3.
+The dialogue-level reward is uniformly attributed to each user utterance, defined as $r_{i,j}^{cc}=\text{Sim}(p_{i},P_{extractor}(d_{i}^{\prime}))$.
+To prevent reward hacking, the AI detection reward is included as an auxiliary component.
+The final reward, defined in Eq.
+5, is utilized to optimize profile recall via Proximal Policy Optimization (PPO)Schulman et al.
+(2017).
 $r_{i,j}=\lambda r^{cc}_{i,j}+(1-\lambda)r^{ai\_detect}_{i,j}$ (5)
-where $r^{ai\_detect}_{i,j}=\text{AI\_detect}(u_{i,j})$ and $\lambda=0.8$ is used to emphasize the importance of cycle consistency.
-The AI detection model Yang et al.
-(2024) and profile generator are all fine-tuned based on our train dataset, details in Appendix B.1.
-4.4 Diverse Profile Sampling
-To generate diverse and naturalistic user profiles that reflect real-world distributions, we propose a Density Profile Sampler that preserves distribution characteristics while ensuring coverage of underrepresented cases.
-Profiles are first embedded into the semantic space using SimCSE Gao et al.
-(2021), then projected with UMAP McInnes et al.
-(2018) to maintain density relationships.
-We then estimate the underlying profile distribution using Gaussian Kernel Density Estimation (GKDE).
-This enables direct sampling of realistic profiles based on probability density, while generating virtual profiles during random sampling by combining OF and SC descriptions from nearest neighbors.
+where $r^{ai\_detect}_{i,j}=\text{AI\_detect}(u_{i,j}^{\prime})$, and $\lambda=0.8$ prioritizes cycle consistency.
+The AI_detect refers to a binary AI detection model Yang et al.
+(2024) that predicts the probability of an utterance being AI-generated.
+Both the AI detection model and profile generator are fine-tuned on our training dataset, with details provided in Appendix B.1.
 5 Experiments
-We evaluate user simulators’ authenticity and consistency at both the utterance and conversation levels while assessing diversity through the difference between our simulated and real user distributions.
+We evaluate user simulators on authenticity, consistency, and multi-turn continuity at both utterance and conversation levels, while measuring diversity by comparing the dialogue distributions of simulated and real users.
 5.1 Datasets
 We select the popular LMSYS-Chat-1M Zheng et al.
-(2023) as our data source for the experiment, which contains one million human-LLM conversations.
-Following previous work Kong et al.
-(2024), we filter the samples for non-English language, toxicity, and redundancy and obtain a complete 94,874 samples (87,882/4,626/2,366 for Training/Validation/Test datasets).
-Then, we use a GPT-4o-based profile extractor described in Section 4.1 to preprocess them with annotating user profiles for each conversation and construct them into LMSYS-USP.
-Detailed preprocessing are in Appendix A.1.
+(2023), which contains one million human-LLM conversations.
+Following prior work Kong et al.
+(2024), we filter out non-English, toxic, and redundant samples, resulting in 94,874 conversations (87,882 for training, 4,626 for validation, and 2,366 for testing).
+Each conversation is then annotated with user profiles using the GPT-4o-based extractor described in Section 4.1, forming the LMSYS-USP dataset.
+Detailed preprocessing steps are provided in Appendix A.1.
 Dataset Profile Source OF SC
@@
 ConvAI2 Human 25.69 8.70 12.64 9.97 16.94 -
-Table 2: Automatic evaluation results of profile quality across different datasets.
-We used DPC and Val.Score to automatically evaluate the quality of extracted user profiles on the test set of LMSYS-USP, plus Persona-Chat Zhang et al.
-(2018) and ConvAI2  [^2] We use the human-to-bot dataset in https://huggingface.co/datasets/convai-challenge/conv_ai_2 Dinan et al.
-(2019) with manually annotated dialogue datasets of other existing predefined profiles.
-Table 2 shows that the extracted profile can achieve over 84% DPC and even the distill-llama3 is close to that of GPT-4o, demonstrating the effectiveness of our annotation.
-Additionally, we select 100 samples for manual evaluation and it shows over 4/5 scores for the quality of the generated profiles (See Appendix B.4 for further details).
+Table 2: Automated evaluation of profile quality across datasets.
+Avg DP.P # Fact denotes the average number of atomic facts per user profile, while Avg DP.R # Fact represents the average number of user utterances per dialogue.
+Note that human-annotated profiles in PersonaChat and ConvAI2 contain no subjective characteristics.
+We use DPC and Val.Score to automatically evaluate the quality of extracted user profiles on 100 randomly selected samples from the LMSYS-USP test set, as well as on 100 samples each from Persona-Chat Zhang et al.
+(2018) and ConvAI2111We use the human-to-bot dataset from https://huggingface.co/datasets/convai-challenge/conv_ai_2 Dinan et al.
+(2019), which include manually annotated predefined profiles.
+As shown in Table 2, the extracted profiles achieve over 84% DPC, with even distill-llama3 results comparable to GPT-4o, demonstrating the effectiveness of our annotation method.
+Manual evaluation further confirms profile quality, with average scores exceeding 4 out of 5 (see Appendix B.3 for details).
+5.2 Configurations
+We train USP based on LLaMA-3-8B-Base AI@Meta (2024) model.
+The conditional SFT is conducted on the training dataset using 4 A100 40GB GPUs, with full fine-tuning over 3 epochs at a learning rate of 5e-5 and max length set to 4096, taking approximately two days.
+Our diverse profile sampler then randomly selects 1,000 samples from the training set for virtual user sampling, combining objective facts and subjective characteristics to generate about 1 million profiles.
+From these, we select the 5,000 profiles least similar to the training dataset for the RLCC phase.
+RLCC training uses two H20 96GB GPUs for 5 days, utilizing a KL coefficient of 0.01, a learning rate of 5e-7, and training for 1 epoch.
 5.3 Baseline Models
-(1) User Simulator without User Profile: This includes the untrained DialogueGPT(4o), where GPT-4o relies solely on context to predict the next user utterance, and PlatoLM Kong et al.
-(2024), which is fine-tuned on our training dataset using LLaMA-3-8B and can be considered equivalent to our approach without incorporating profile.
-(2) User Simulator Guided by the User Profile: We adopt ProfileGPT(4o) and ProfileGPT(llama), which utilizes GPT-4o and LLAMA-3-8B-Instruct AI@Meta (2024) as the role-play backbone with our constructed profiles.
-Additionally, we include CharacterGLM Zhou et al.
-(2024a), which performs role-playing with any given profile, and CharacterLLM Shao et al.
-(2023), which role-played on several famous celebrities.
-The detailed setup and prompts are provided in Appendix B.3.
+(1) User Simulator without User Profile: This includes untrained GPT-4o (User w/o Profile), which use GPT-4o to predict user utterances based solely on context, and PlatoLM Kong et al.
+(2024), a baseline fully fine-tuned on our dataset using LLaMA-3-8B-Base, representing a profile-agnostic approach.
+(2) User Simulator with User Profile: We employ GPT-4o (User w/ Profile) and LLaMA3 (User w/ Profile) leveraging GPT-4o and LLaMA-3-8B-Instruct AI@Meta (2024) as profile-conditioned role-playing agents, alongside CharacterGLM Zhou et al.
+(2024a), a flexible profile-based baseline, and CharacterLLM Shao et al.
+(2023), designed to emulate public figures.
 5.4 Metrics
-Authenticity: We use SimCSE Gao et al.
-(2021) to compute semantic similarity (Sem-Sim) and style embeddings Wegmann et al.
-(2022) to compute style similarity (Style-Sim) for evaluating $D_{\text{utt}}(u_{ij},u^{\prime}_{ij})$ and $D_{\text{dia}}(d_{i},d_{i}^{\prime})$.
-We also employ Author Verification Accuracy (AVA) to assess stylistic consistency by measuring whether paired sentences share authorship based on similarity thresholds Wegmann et al.
-(2022).
-For multi-turn evaluation, we compute dialogue-level distances by concatenating each user’s utterances.
-Consistency: We evaluate dialogue-level consistency between dialogues and profiles using reverse metrics: r-DP.P and r-DP.R, calculated as $Fact.Con(P_{i},d_{i}^{\prime})$ and $Fact.Con(d_{i}^{\prime},P_{i})$ respectively, along with their harmonic mean r-DPC, from a profile-centric perspective.
-For utterance-level evaluation, we compute the mean DP.P score across individual utterances.
-Additionally, we incorporate Persona Coverage (P.Cover) Song et al.
-(2019) to assess keyword-level consistency.
-The Subjective Characteristic Score (SC.Score), measures the reflection of subjective traits, assessed by GPT-4o with prompt in Appendix D.
-Diversity: We compute the Absolute Difference Value (ADV), which represents the Euclidean distance between the PCA-reduced embeddings of generated and target dialogues, to evaluate the discrepancy between the distribution of the reconstructed dialogues and the original dialogues.
-Additionally, we assess multi-turn dialogue continuity using the early stop rate (ESR), which flags premature endings caused by repetitive responses or repeated gratitude expressions over three turns.
+Authenticity: We evaluate semantic and stylistic similarity using SimCSE Gao et al.
+(2021) and style embeddings Wegmann et al.
+(2022), respectively, to compute $D_{\text{utt}}(u_{i,j},u^{\prime}_{i,j})$ and $D_{\text{dia}}(d_{i},d_{i}^{\prime})$.
+To assess stylistic consistency, we report Author Verification Accuracy (AVA) Wegmann et al.
+(2022), which measures whether sentence pairs are attributed to the same author based on similarity thresholds.
+Dialogue-level distances are computed by concatenating all user utterances.
+[…]
+Diversity: We measure the Absolute Difference Value (ADV), defined as the Euclidean distance between PCA-reduced embeddings of generated and target dialogues, to quantify the distributional discrepancy between simulated and target dialogues.
+Continuity: Multi-turn dialogue continuity ability is evaluated via the early stop rate (ESR), which detects premature endings triggered by repetitive responses or repeated expressions of gratitude across three turns.
 5.5 Results
@@
 […]
-w/o Profile DialogueGPT(4o) 40.24 13.75 11.28
+w/o Profile GPT-4o (User w/o Profile) 40.24 13.75 11.28
 PlatoLM 39.37 43.11 40.29
-With Profile Character_LLM 37.54 18.88 15.03 54.77 66.62 2.43
+w/ Profile Character_LLM 37.54 18.88 15.03 54.77 66.62 2.43
 Character_GLM 38.51 22.28 18.17 68.72 57.72 2.95
-ProfileGPT(llama) 39.82 14.88 13.47 82.19 72.29 3.92
-ProfileGPT(4o) 41.66 5.74 9.87 92.73 73.34 4.71
-USP w/o RLCC 54.25 46.57 43.61 71.30 71.56 3.36
+LLaMA3 (User w/ Profile) 39.82 14.88 13.47 82.19 72.29 3.92
+GPT-4o (User w/ Profile) 41.66 5.74 9.87 92.73 73.34 4.71
+USP (w/o RLCC) 54.25 46.57 43.61 71.30 71.56 3.36
 USP 53.38 46.60 43.35 72.61 71.23 3.39
-Table 3: Utterance-level performance comparison of different models on authenticity and consistency metrics.
+Table 3: Utterance-level performance comparison of different user simulator.
+In utterance-level evaluation, we assess the quality of single-turn responses generated by different user simulators given the golden context.
+As shown in Table 3, USP outperforms all baselines in authenticity, as measured by both semantic similarity (Sem-Sim: 53.38) and stylistic similarity (Style-Sim: 46.60).
+This shows the effectiveness of our implicit profile-based approach for user-LLM dialogue reconstruction, especially compared to non-profile baselines like PlatoLM.
+While dedicated role-playing models (e.g., GPT-4o (User w/ Profile)) achieve higher consistency scores (r-DP.P) due to direct profile keyword copying with high P.Cover (73.34), our USP strikes a better balance between authenticity and consistency, as shown by the intuitive examples in Section 6.3.
+5.5.2 Conversation-Level Evaluation
 Model Type Model Continuity Authenticity Consistency
 […]
-w/o Profile DialogueGPT(4o) 35 48.91 14.21 10.58
+w/o Profile GPT-4o (User w/o Profile) 35 48.91 14.21 10.58
 PlatoLM 18 43.24 32.43 31.60
-With Profile Character_LLM 52 23.37 7.13 4.69 25.48 6.43 10.27 21.49 2.82
+w/ Profile Character_LLM 52 23.37 7.13 4.69 25.48 6.43 10.27 21.49 2.82
 Character_GLM 44 40.19 10.86 12.67 39.51 29.61 33.85 42.75 3.64
-ProfileGPT(llama) 31 46.84 10.58 11.63 67.09 29.98 41.44 47.72 4.19
-ProfileGPT(4o) 32 48.87 10.15 11.26 76.59 43.72 55.66 51.02 4.56
-USP w/o RLCC 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
+LLaMA3 (User w/ Profile) 31 46.84 10.58 11.63 67.09 29.98 41.44 47.72 4.19
+GPT-4o (User w/ Profile) 32 48.87 10.15 11.26 76.59 43.72 55.66 51.02 4.56
+USP (w/o RLCC) 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
 USP 10 65.39 46.23 38.77 56.24 74.38 64.05 44.08 3.35
-Table 4: Conversation-level performance comparison of different models on authenticity and consistency metrics.
-In the utterance-level evaluation, we evaluate the quality of a single-turn response generated by the testing models in a given context.
-As shown in Table  3, USP outperforms all baselines in terms of authenticity, with 53.38 and 46.60, as measured by both semantic (Sem-Sim) and stylistic (Style-Sim) similarity metrics.
-This highlights the effectiveness of our implicit profile-assisted approach for user-LLM dialogue reconstruction, particularly when compared to context-only models like PlatoLM.
-Although dedicated role-playing models (ProfileGPT variants) achieve higher consistency scores (r-DP.P), this can be attributed to their direct profile text copying.
-USP maintains comparable overall performance while striking a better balance between authenticity and consistency.
-5.4.2 Conversation-Level Evaluation
-In the conversation-level evaluation, we assess the performance of testing models to chat with GPT-4o in multi-turns, providing the profile or the first turn of the reference dialogue according to their needs.
+Table 4: Conversation-level performance comparison of different user simulators.
+In the conversation-level evaluation, we assess the quality of multi-turn dialogues generated by different user simulators interacting with GPT-4o, each provided with either a given profile or the first turn of a reference dialogue.
 As shown in Table 4, USP outperforms baseline models in authenticity, consistency, and continuity.
-With the lowest ESR(10), it demonstrates superior dialogue continuity.
-Notably, USP’s advantage in authenticity is more pronounced in conversation-level scenarios than sentence-level predictions.
-In terms of consistency, USP demonstrates exceptional performance in r-DP.R metrics and achieves significantly higher r-DPC scores for overall profile dialogue consistency.
-This superior performance, particularly when compared to role-playing models such as ProfileGPT(4o) and ProfileGPT(llama) which show high P.Cover, suggests that RLCC effectively captures abstract profile characteristics beyond merely matching surface-level keywords.
-See Appendix C.2 for further case study analysis.
+With the lowest ESR (10), USP ensures superior dialogue continuity, avoiding issues like repetitive generation and reciprocal appreciation loops seen in baselines.
+Its advantage in authenticity is especially evident in multi-turn scenarios, compared to sentence-level evaluations.
+In terms of consistency, USP excels with a high r-DP.R (74.38) and significantly better r-DPC (64.05), demonstrating strong conditional generation consistency.
+Unlike role-playing models such as GPT-4o (User w/ Profile) and LLaMA3 (User w/ Profile), which show high P.Cover and r-DP.P but lower overall profile dialogue consistency, USP demonstrates a deeper and more comprehensive understanding of user behavior, moving beyond surface-level keyword matching to deliver a more vivid user simulation.
 5.5.3 Human Evaluation
-We also randomly selected 100 samples with 8 evaluators to conduct conversation-level human evaluations considering authenticity and consistency.
-Authenticity was assessed through Style, Semantics, and Quality, while consistency covered Accuracy, Completeness, and Quality.
-Full evaluation details are provided in Appendix B.4.
-Table 5 demonstrates USP’s superior performance in both authenticity and consistency metrics.
-Our USP significantly outperforms ProfileGPT (4o) in terms of authenticity (74 vs.
+We randomly selected 100 samples from the test set and engaged 8 evaluators to assess conversations on authenticity and consistency.
+Authenticity was evaluated based on Style, Semantics, and Quality, while consistency focused on Accuracy, Completeness, and Quality.
+Detailed criteria are in Appendix B.3.
+Table 5 shows USP’s clear superiority in both authenticity and consistency.
+USP outperforms GPT-4o (User w/ Profile) in authenticity (74 vs.
 13) and consistency (61 vs.
-35) in manual evaluation.
-USP is superior to PlatoLM trained on the same dataset in terms of authenticity, which demonstrates the usefulness of implicit profile modeling.
-Thanks to the RLCC module, our USP model has significantly improved consistency (43 vs.
-30) by aligning user profiles.
+35).
+It also surpasses PlatoLM trained on the same data in authenticity, demonstrating the advantage of implicit profile modeling.
+The larger gap in consistency (43 vs.
+30) compared to authenticity (37 vs.
+31) between USP and USP (w/o RLCC) highlights RLCC’s key role in aligning profiles with dialogues.
 Baseline Metrics (USP win/tie/loss)
@@
 ($\kappa$=0.548) ($\kappa$=0.561)
-ProfileGPT(4o) 74/13/13 61/4/35
+GPT-4o (User w/ Profile) 74/13/13 61/4/35
 PlatoLM 55/12/33 -
-USP w/o RLCC 37/32/31 43/27/30
-Table 5: Human evaluation results comparing baselines with USP on authenticity and consistency.
+USP (w/o RLCC) 37/32/31 43/27/30
+Table 5: Human evaluation of USP win rates over baselines in terms of authenticity and consistency.
+$\kappa$ denotes the within-group kappa coefficient.
+Note that PlatoLM, as a non-profile baseline, contain no consistency results.
 5.5.4 Diversity Sampling Evaluation
+Figure 3: Cumulative distribution of ADV performance comparison across different simulators, with the red cross indicating that USP (w/o RLCC) has 60% of its samples with ADV below 5%.
+Figure 3 compares ADV between target and generated dialogues across simulators and percentiles.
+The USP series consistently demonstrate lower ADV, even in extreme cases, with 60% of samples achieving ADV below 5% (marked by a red cross), compared to baselines(e.g.
+PlatoLM, GPT-4o (User w/ Profile)) at 10% or higher.
+This reflects the USP series’ superior ability to generate dialogues that closely align with target conversations, effectively preserving the diversity distribution of user characteristics.
+We also show that our sampling strategy outperforms random sampling by effectively capturing diverse representatives (majority and minority) in Appendix C.1.
+6 Analysis
+6.1 Ablation Study
 Model Configuration Continuity Authenticity Consistency
 […]
-USP w/o RLCC 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
+USAP (w/o RLCC) 17 64.22 40.96 35.55 65.36 51.29 57.47 50.35 3.21
+USP (w/o RLCC) 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
 USP (5:5) 14 66.28 41.22 37.03 52.23 71.59 60.39 43.58 3.55
@@
 USP (9:1) 12 66.91 38.87 33.62 58.36 70.62 63.90 46.75 3.33
-Table 6: Ablation study of hyperparameters in RLCC.
-Figure 3 shows the absolute difference value between target dialogue and generated dialogue by various models across different percentiles.
-From the results, we observe that USP and USP w/o RLCC consistently achieve the smallest ADV across all percentiles, indicating the dialogues they generated that closely match the target conversations.
-For example, marked by the red cross, PlatoLM has 60% of sample ADV below 15%, while USP has only 5% or less ADV.
-It demonstrates its stronger capability in preserving the semantic characteristics of the original dialogues.
-The uniformly lower curves of USP and USP w/o RLCC compared to other baselines (PlatoLM, ProfileGPT(4o), and DialogueGPT(4o)) suggest that our approach generates dialogues that are more faithful to the target conversations across typical and extreme cases.
-We also analyze and demonstrate that our user simulator can sample different representatives (majority and minority) of users compared to random sampling in Appendix C.1 and apply it to downstream applications in Appendix C.3.
-Figure 3: Cumulative distribution of ADV performance comparison across different models.
-6 Analysis
-6.1 Ablation Study
-To evaluate the relative importance of RLCC’s two rewards, we tested different values of $\lambda$ in Equation 6, denoted as USP($\lambda$: $1-\lambda$).
-As shown in Table 6, $\lambda$ = 0.8 provides the optimal balance between maintaining model capabilities and enhancing dialogue consistency.
-Higher values of $\lambda$ (0.9) compromise speaking style authenticity without improving r-DPC, leading to superficial profile matching, as shown by increased P.Cover scores.
-Conversely, $\lambda$ = 0.5 achieves authentic style features but lacks sufficient consistency emphasis, resulting in stagnant performance across capabilities.
+Table 6: Ablation study of our USP framework.
+We evaluate the effectiveness of the polishing step in our two-stage profile construction pipeline, which converts attributes into natural language descriptions, by comparing it to a baseline (USAP w/o RLCC) that uses only attributes without polishing.
+As shown in the first two rows of Table 6, the polishing step enhances generalization, improving performance across most metrics of Continuity, Authenticity, and Consistency.
+In contrast, relying solely on attributes leads USAP (w/o RLCC) to excessively replicate profile descriptions, resulting in a high P.Cover score (50.35) due to attributes appearing directly in the dialogue.
+We also assess the relative importance of RLCC’s two rewards by testing different $\lambda$ values in Equation 5, denoted as USP($\lambda$: $1-\lambda$).
+Table 6 shows that $\lambda=0.8$ optimally balances model capabilities and dialogue consistency.
+Higher $\lambda$ (0.9) sacrifices speaking style authenticity without improving r-DPC, increasing P.Cover and indicating superficial profile matching.
+Conversely, $\lambda=0.5$ preserves authentic style but lacks sufficient consistency, resulting in stagnant performance.
 6.2 Applications: Dynamic Multi-turn Evaluation For LLMs
-One application of our simulator is to fill the gap in the current dynamic multi-turn evaluation of large models.
-The user simulator can simulate different user groups to dynamically interact with the tested model in multiple rounds and reveal their specific defects, as shown in Table 7.
+One application of our simulator is addressing the gap in dynamic multi-turn evaluation of LLMs.
+While current automatic evaluations rely on static preset questions Zheng et al.
+(2024); Bai et al.
+(2024), our user simulator can dynamically interact with the LLM over multiple rounds, adjusting based on response quality and given user traits.
+We generated 300 diverse user profiles using our sampler: 100 highest-probability (majority), 100 lowest-probability (minority), and 100 random synthetic (virtual) profiles.
+Using these profiles, USP engages in multi-turn dialogues with the LLM.
+Evaluation results, based on MT-Bench (Zheng et al., 2024) and presented in Table 7, show that dynamic multi-turn evaluation aligns closely with average rankings on LiveBench White et al.
+(2024) and Chatbot-Arena Chiang et al.
+(2024), while revealing fine-grained weaknesses of different LLMs when interacting with specific user groups.
+Detailed analysis is in Appendix C.3.
 Model Setup Sampling Strategy Avg. Ranking in LiveBench/
@@
 Claude-Haiku 4.88 5.42 5.43 5.24 5
-Table 7: Response model performance comparison over different target groups.
-We simulated 300 diverse user profiles by USP, comprising 100 highest-probability profiles(the majority), 100 lowest-probability profiles(the minority) based on estimated density, and 100 random synthetic profiles(the virtuality), using the sampler mentioned in Section 4.4.
-Then, we leverage USP using these profiles to chat with LLM in multi-turns and evaluate them, followed by MT-Bench (Zheng et al., 2024).
-The experimental results indicate that our user simulator is effective and consistent with the average rankings on the latest Livebench White et al.
-(2024) and Chatbot-Arena Chiang et al.
-(2024).
-In addition, it can also be seen that GPT-4o is better at catering to minority groups than other models, indicating its superior robustness.
-A more detailed analysis can be found in the appendix C.3.
+Table 7: LLM performance across user groups.
+6.3 Case Study
+Table 8 shows that the role-playing baseline GPT-4o (w/ Profile) often copies abstract profile traits verbatim to assert user identity.
+In contrast, our USP conveys these traits more naturally by transforming abstract concepts into concrete and coherent expressions.
+For instance, when the profile states “being a father of two,” GPT-4o (w/ Profile) repeats it directly, while USP implicitly reflects this by mentioning a “son” and a “daughter” in later turns.
+Similarly, rather than restating “likes Italian food,” USP refers to a specific dish like pasta.
+These examples illustrate that USP better mimics human behavior by expressing abstract, high-level traits both directly and subtly in a natural manner, likely contributing to its stronger human preference (see Table 5).
+Additional case studies are provided in Appendix C.2.
+(Profile) As a father of two, your strong sense of family responsibility is evident… Your interest in vegan Italian-inspired snack dishes
+(1st turn) GPT-4o (w/ Profile): As a father of two who has a keen interest in …
+(2nd turn) GPT-4o (w/ Profile): Lately, I’ve been exploring new vegan Italian-inspired snack dishes
+(4th turn) USP:My son wants pasta but he doesn’t like tomato sauce…
+(5th turn)my daughter says she does not feel well… She prefers vegan food in Italian style
+Table 8: Case study comparing outputs from USP and the GPT-4o (w/ Profile) baseline for a sample profile.
+Bold highlights keywords explicitly copied by GPT-4o, while bold italic marks USP’s implicit or fuzzy matches.
 7 Conclusion
-In this work, we propose a novel user simulator with implicit profiles that excels in authenticity, consistency, and diversity.
-Based on this, we introduce the USP framework, which integrates extracted user profiles into the user simulator by conditional fine-tuning and reinforcement learning with cycle consistency.
-Our experimental results, validated by both automatic metrics and human evaluations, show that USP significantly outperforms role-playing simulators (e.g., GPT-4o) and direct simulation approaches (e.g., PlatoLM) in authenticity while achieving comparable consistency at both the sentence and conversation levels.
-Furthermore, through a dynamic evaluation across various LLMs chatting with diverse demographic groups, we demonstrate USP’s effectiveness in real-world applications.
+In this work, we introduce the USP framework, which integrates extracted user profiles into the user simulator by conditional SFT and RLCC.
+Our experimental results, validated by both automatic metrics and human evaluations, show that USP significantly outperforms role-playing simulators (e.g., GPT-4o (User w/o Profile)) and direct simulation approaches (e.g., PlatoLM) in authenticity and diversity while achieving comparable consistency at both the sentence and conversation levels.
+Additionally, dynamic evaluations with various LLMs across diverse demographic groups highlight USP’s effectiveness in real-world scenarios.
+Nonetheless, a gap remains compared to real human behavior, and our future work will explore finer-grained control and multimodal simulation.
 Limitations
-We acknowledge the following limitations: 1) Applicability Across Different Scenarios: We conduct the experiments on a single dataset, and there has been limited validation across multiple datasets to assess the generalizability of the results.
-2) Cultural and Linguistic Scope: We focus on English dialogues in this paper, which may limit the applicability of USP to other linguistic and cultural contexts.
-Ethics Statement
-Although LMSYS-1M has undergone extensive data cleaning and ethical checks, the dataset may still contain sensitive or harmful content, reflecting violent, explicit, or discriminatory traits in certain dialogue.
-This could result in USP generating unsafe dialogues.
-We strongly advise against including sensitive terms in profiles when using USP, as this may lead to extreme behavior in both the USP and the response model.
+We acknowledge the following limitations: 1) Scenario Applicability: Experiments were conducted on a single dataset, with minimal validation across others to confirm broader applicability.
+2) Linguistic and Cultural Scope: Our focus on English dialogues may limit the applicability of USP to other languages and cultural contexts.
+Acknowledgments
+This research is supported by the project of Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No.
+JCYJ20220818103001002), Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No.
+ZDSYS20230626091302006), Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing, SRIBD Innovation Fund (Grant No.
+K00120240006), and Program for Guangdong Introducing Innovative and Entrepreneurial Teams, Grant No.
+2023ZT10X044.
 Appendix A Dataset Construction
@@
 A.2 Profile Dataset
-As discussed in Section 4.1, we categorize all attributes into three types: scene-consistent attributes, scene-related attributes, and deep intrinsic characteristics.
-For scene-consistent attributes, we use the prompt shown in Figure 13, with each metric definition following the guidelines outlined in Cheng et al.
-(2024b).
-For scene-related attributes, we use the prompt in Figure 14, and for deep intrinsic characteristics, we refer to Figure 12.
-The definition of the Big Five Traits scores follows Gosling et al.
+As detailed in Section 4.1, we classify attributes into three types: scene-consistent, scene-related, and deep intrinsic characteristics.
+For each, we use specific prompts (Figures 11, 12, and 13), with metric definitions based on Cheng et al.
+(2024b) and Big Five traits per Gosling et al.
 (2003).
-Next, we concatenate the attributes, remove invalid values (e.g., null or meaningless values), and shuffle the order to eliminate any positional bias in the generated profiles.
-The attributes, which encompass the three aforementioned aspects, are then rephrased using GPT-4o, with the prompt shown in Figure 15.
-As a result, we obtain automatically labeled profiles for each data entry.
-The length statistics are shown in Table 8.
-Dataset Train Val Test Profile
-LMSYS-USP 1,149 1,295 1,438 231
-Table 8: Average token length of LMSYS-USP dataset.
-Furthermore, we calculated the frequency of occurrence for each attribute value (i.e., the average number of different attribute values per sample) to assess the prevalence of each attribute.
-The statistics for the objective facts can be found in Figure 4, while for the subjective characteristics, we focused on whether the Big Five Traits were significantly exhibited.
-Specifically, we only consider traits with high or low scores, while moderate scores are viewed as the average representation of human behavior Moon et al.
-(2024) and are not included in the subsequent profiles.
+We then concatenate these attributes, remove invalid entries, and shuffle their order to prevent positional bias.
+The combined attributes are rephrased using GPT-4o with the prompt in Figure 14, producing automatically labeled profiles.
+The LMSYS-USP dataset averages 1,149 tokens in training, 1,295 in validation, 1,438 in testing, and 231 tokens per profile.
+We also measured the frequency of each attribute value, defined as the average number of distinct values per sample, to assess attribute prevalence.
+Statistics for objective facts are shown in Figure 4.
+For subjective traits, we focused on the Big Five traits only when scores were significantly high or low, excluding moderate scores as they reflect average human behavior Moon et al.
+(2024) and were omitted from the subsequent polishing step.
 Attribute High Rate (%) Low Rate (%)
@@
 Table 9: Summary of extracted subjective attribute statistics.
-Figure 4: Frequency of occurrence of values across different attributes of objective facts in the attribute extraction process.
+Figure 4: Frequency of values for each attribute of objective facts in the attribute extraction process.
 A.3 Resource Consumption in Implementation
-During the data construction process, each attribute extraction requires approximately \$0.003 using the GPT API.
-Since each sample requires three extractions, the cost per sample is approximately \$0.01.
-With a total of around 94,000 samples, the cost for attribute extraction amounts to approximately \$940.
-Additionally, rewriting the attributes into profiles incurs a cost of around \$0.05 per sample.
-Therefore, the total cost for constructing the dataset is approximately \$1,400.
+Attribute extraction using the GPT-4o API costs  \$0.003 per attribute type, or  \$0.01 per sample for three types.
+For 94,000 samples, the extraction costs  \$940.
+Rewriting attributes into profiles adds  \$0.05 per sample, resulting in a total dataset construction cost of  \$1,400.
 Appendix B Implement Detail
 B.1 Trainable Model Setup
-We use the llama3-base model as the base architecture.
-Directly training an instruct model with user simulation contradicts the inherent task requirements during its SFT phase, making convergence difficult.
-We first perform conditional SFT on the training dataset, followed by our diverse profile sampler, which randomly selects 1,000 samples from the training set for virtual user sampling.
-Specifically, we combine objective facts and subjective descriptions from different profiles and generate approximately 1 million profiles.
-From these, we select the 5,000 profiles with the lowest similarity to the training dataset and use them for RLCC phrase.
-The conditional SFT is conducted using four A100 40GB GPUs for full fine-tuning with epoch set to 3, taking about two days.
-The RLCC phase is trained using two H20 96GB GPUs over the course of five days.
-For the PlatoLM, we also use llama3-base as the base architecture.
-The system prompt used is: “A chat between a curious human and an artificial intelligence assistant.
-The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.” We fine-tune the model using four A100 40GB GPUs with epoch set to 3, which takes approximately two days.
-For the AI detection model, we follow Cheng et al.
-and use Longformer Beltagy et al.
-(2020) to train on our dataset.
-Since our dataset naturally distinguishes between AI and human-generated text, we label user utterances as human and assistant utterances as AI.
-We trained for 3 epochs using dual 3090 GPUs, taking three days to complete
-For the profile generator model, we utilized LLaMA3-Instruct AI@Meta (2024) as the backbone and fine-tuned it on our carefully curated profile dataset.
-This approach effectively distilled the two-stage profile generation capabilities of GPT-4o.
-Training was conducted on four A100 40GB GPUs for 3 epochs, taking two days to complete.
-B.2 Baseline Model Setup
-Table 10 presents the experimental setup for the baseline models.
-For models based on GPT, we use the corresponding APIs with default settings for inference.
-For other models, the experiments are conducted on a single NVIDIA RTX A100 GPU with a batch size of 2 and a repetition penalty of 1.0.
-An exception is made for CharacterGLM, where we utilize its chat function with a repetition penalty of 1.6 and set the number of beams to 3.
-Model Name Backbone System Prompt
-USP LLAMA3-8B
-You are engaging in a conversation with an AI assistant. your profile is:
-{profile}.
-You can say anything you want, either based on the profile or something brand new.
-DialogueGPT(4o) gpt-4o-2024-08-06 -
-PlatoLM LLAMA3-8B
-A chat between a curious human and an artificial intelligence assistant.
-The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.
-ProfileGPT(4o) gpt-4o-2024-08-06
-You are engaging in a conversation with an AI assistant. your profile is:
-{profile}.
-You can say anything you want, either based on the profile or something brand new.
-CharacterGLM ChatGLM-6B
-以下是一段User和AI assistant之间的对话。
-关于User的信息:{profile}
-关于AI assistant的信息:GPT-4o
-Character-LLM-Socrates-7b LLAMA-7B
-I want you to act like the person described in the profile below: {profile}.
-I want you to respond and answer like the person, using the tone, manner, and vocabulary the person would use.
-You must know all of the knowledge of the person.
-The status of you is as follows:
-Status: You are engaging in a conversation with an AI assistant.
-The interactions are as follows:
-Table 10: Baseline model setup.
-B.3 Seed conversation design
-As there are two types of simulators, one being a response model based on role-playing, which cannot proactively initiate conversations, we first embed the corresponding profile into the system prompt.
-Then, by using the query "What will you say to start the conversation?" we guide the model to simulate the user’s input over the test dataset.
-Finally, the simulated user utterance is passed to the response model for interactive generation.
-The second type of simulator can directly generate reasonable user utterances without requiring additional guidance after setting up the system prompt.
-Specific configurations of the system prompts for different baseline models are outlined in Table  10.
-B.4 Human Evaluation
-B.4.1 Profile Evaluation
-We employed two annotators to rate the extracted profiles on a scale of 1 to 5 based on the given dialogues, assessing the accuracy and completeness of objective facts, reasonableness of subjective descriptions, and overall naturalness.
-For accuracy, we emphasized precision without hallucinations; for completeness, we focused on the absence of significant omissions; and for reasonableness, we prioritized rational, well-justified descriptions, avoiding exaggeration or bias.
-The results in Table 11 show that annotators achieved moderate to high agreement on profile assessment.
-The evaluation interface is illustrated in Figure 6.
+For PlatoLM, we base it on the LaMA-3-8B-Base architecture.
+Following Kong et al.
+(2024), the system prompt is: "A chat between a curious human and an artificial intelligence assistant.
+The human can ask follow-up or new questions without prior context." We fine-tune using four A100 40GB GPUs for 3 epochs, taking about two days.
+The AI detection model uses Longformer Beltagy et al.
+(2020) trained on our dataset per Cheng et al.
+.
+User utterances are labeled as human, and assistant utterances as AI.
+Training runs for 3 epochs on dual RTX 3090 GPUs, taking three days.
+The profile generator is fine-tuned from LLaMA-3-8B-Instruct AI@Meta (2024) on our curated profile dataset, effectively distilling GPT-4o’s two-stage profile generation.
+Training uses four A100 40GB GPUs for 3 epochs, lasting two days.
+B.2 Train-Free Model Setup
+We use two simulator types: (1) A response model, e.g., GPT-4o (User with Profile), role-playing as a user simulator, unable to initiate conversations.
+The user profile is embedded in the system prompt, and the initial query is generated by asking, "What will you say to start the conversation?" to obtain the user’s opening query.
+(2) A user simulator, e.g., PlatoLM, directly generates user utterances from a seed prompt without additional steps.
+B.3 Human Evaluation
+All annotators we recruited were based on two criteria: (1) an IELTS score of 6.5 or higher for sufficient English proficiency, and (2) a Computer Science background with research experience or foundational knowledge in dialogue systems.
+B.3.1 Profile Evaluation
+Two annotators (one undergraduate, one master’s student) rated extracted profiles on a 1–5 scale based on dialogues, assessing: (1) accuracy of objective facts (precision without hallucinations), (2) completeness (no significant omissions), and (3) reasonableness of subjective descriptions (rational, unbiased, justified).
+Results in Table 10 indicate moderate to high annotator agreement.
 Dataset Profile Source Objective Facts Subjective Characters Naturalness
 LMSYS-USP GPT-4o 4.64 4.19 4.66
-Table 11: Human evaluation results for the quality of constructed profiles.
-B.4.2 Dialogue Evaluation
-We recruited eight annotators to assess conversation-level results based on two key criteria: authenticity and consistency.
-For authenticity, annotators were tasked with identifying user utterances within a dialogue that most closely resemble those in a reference dialogue, evaluating them across three dimensions: Style, Semantics, and Quality.
-For consistency, annotators were provided only with the user profile and asked to select the user utterances that best align with the profile, considering Accuracy, Completeness, and Quality.The definitions follow prior work Cheng et al.
-(2024a, b), with detailed guidelines illustrated in Figure 16 and Figure 17.
-To minimize position bias and the potential influence of prior exposure, each annotator was randomly assigned pairs of dialogues, with the order of options also randomized.
-The evaluation interfaces for authenticity and consistency assessments are presented in Figure 7 and Figure 8, respectively.
+Table 10: Human evaluation results for profile quality across three aspects: Objective Facts, Subjective Characters, and Naturalness, with 1-point agreement rates of 89.2%, 74.3%, and 88.4% respectively.
+B.3.2 Dialogue Evaluation
+We recruited eight annotators—comprising two undergraduates, five postgraduates, and one postdoctoral researcher—to evaluate conversation-level results.
+This diverse academic representation ensured a broad range of expertise.
+Annotators assessed dialogues using two key criteria: authenticity and consistency.
+For authenticity, they compared user utterances against a reference dialogue, focusing on style, semantics, and quality.
+For consistency, annotators evaluated user utterances solely based on the provided user profile, considering accuracy, completeness, and quality.
+These definitions align with prior work Cheng et al.
+(2024a, b), with detailed guidelines provided in Figure 7 and Figure 8.
+Eight annotators were randomly paired into four groups, each independently evaluating a randomly assigned dialogue sample.
+Each group reviewed 75 examples across three baselines (100 examples each).
+To reduce position bias and prior exposure, dialogue pair assignments and their presentation order were randomized.
 Figure 5: Distribution of different sampling strategies.
-Figure 6: Human evaluation interface for extracted profile evaluation.
-Figure 7: Human evaluation interface for assessing dialogue authenticity.
-Evaluators compare two generated conversations and select which user’s utterances more closely align with those in the reference dialogue.
-Figure 8: Human evaluation interface for assessing dialogue consistency. Evaluators determine which user from two test conversations better aligns with a provided user profile.
 Appendix C Further Analysis
 C.1 Sampling strategy effectiveness
-To evaluate our density sampler’s effectiveness, we use two complementary metrics: Local Density Loss (LDL) Rodriguez and Laio (2014) for structure preservation and Uniformity Loss Wang and Isola (2020) for global coverage.
-Lower LDL values indicate tighter local clustering, confirming preservation of natural profile distribution structures, while lower Uniformity Loss values demonstrate better global coverage while maintaining realistic distributions.
-Guided by GKDE density distribution, we implement two targeted strategies: sampling high-density regions for majority patterns and weighting low-density regions for minority case coverage.
-Figure 5 demonstrates how this approach maintains distribution balance while achieving these objectives.
-Moving right to left on the uniformity loss axis reveals a clear transition from majority to minority sampling patterns.
-Majority samples (blue circles) perform optimally in low LDL regions with high uniformity, preserving key distribution characteristics.
-Minority samples (orange squares) strategically occupy higher density regions with moderate uniformity, capturing diverse distribution patterns.
-The steady progression of sampling percentages in both cases demonstrates controlled sampling behavior, while random sampling (green triangles) shows more scattered patterns, validating our approach’s reliability.
-The overall performance marker (red star) at the intersection confirms our method successfully balances distribution preservation and targeted sampling objectives.
+To evaluate our density sampler, we use two complementary metrics: Local Density Loss (LDL)Rodriguez and Laio (2014) to assess structure preservation, and Uniformity LossWang and Isola (2020) to measure global coverage.
+Lower LDL indicates tighter local clustering, preserving natural profile structures, while lower Uniformity Loss reflects better global coverage with realistic distributions.
+Guided by GKDE density estimates, we apply two strategies: sampling high-density regions to capture majority patterns and weighting low-density regions to cover minority cases.
+Figure 5 shows how this approach balances distribution preservation and targeted sampling.
+Moving along the uniformity loss axis reveals a shift from majority samples (blue circles), which excel in low LDL and high uniformity regions, to minority samples (orange squares), which occupy higher LDL areas with moderate uniformity to capture diversity.
+Sampling percentages progress steadily for both, indicating controlled behavior, while random sampling (green triangles) displays scattered patterns, confirming our method’s reliability.
+The overall performance (red star) highlights a successful balance between preservation and targeted sampling.
 C.2 Case Study
-To evaluate USP’s specific performance in terms of consistency and authenticity, we present two non-cherry-picked dialogues generated through interactive conversations with GPT-4o.
-Figure 10 examines authenticity by comparing against reference dialogues and other baselines, while Figure 11 evaluates profile consistency among profile-based baseline models.
-In terms of authenticity, USP successfully captures both stylistic nuances (such as the consistent use of lowercase "i" and concise questioning patterns) and semantic alignment with target conversations.
-In contrast, PlatoLM deviates from the dialogue flow by the fourth turn, while ProfileGPT(4o) falls into repetitive appreciation patterns.
-Regarding consistency, USP effectively synthesizes both objective profile information and subjective characteristics, demonstrating robust generalization capabilities when handling previously unseen user profiles.
+To evaluate USP’s performance on consistency and authenticity, we present two dialogues generated via interactive conversations with GPT-4o.
+Figure 9 assesses authenticity by comparing USP with reference dialogues and other baselines, while Figure 10 examines profile consistency across profile-based models.
+USP captures stylistic nuances—such as the consistent use of lowercase “i” and concise questioning—and maintains strong semantic alignment with target conversations.
+In contrast, PlatoLM diverges from the dialogue flow by the fourth turn, and GPT-4o (User w/ Profile) falls into repetitive praise.
+On consistency, USP effectively integrates objective profile details and subjective traits, demonstrating strong generalization to unseen user profiles.
 C.3 Downstream analysis
@@
 (c) Performance on the virtual group
-Figure 9: Performance trends of different LLM models across dialogue turns for various demographic groups.
-Our analysis of performance trends across dialogue turns for mainstream LLMs with different demographic groups reveals four key findings, as illustrated in Figure 9: (1) While LLMs demonstrate robust performance with majority demographics, they show notably decreased overall effectiveness when interacting with minority groups, highlighting limitations in personalization capabilities; (2) The models maintain reasonable performance with virtual groups, suggesting effective generalization abilities beyond real-world demographics; (3) Instruction-following capability gradually declines as dialogue turns increase, aligning with observations from previous studies Kwan et al.
+Figure 6: Performance trends of different response models across dialogue turns for various demographic groups.
+Our analysis of performance trends across dialogue turns for mainstream LLMs with different demographic groups reveals four key findings, as illustrated in Figure 6: (1) While LLMs demonstrate robust performance with the majority demographics, they show notably decreased overall effectiveness when interacting with minority groups, highlighting limitations in personalization capabilities; (2) The models maintain reasonable performance with virtual groups, suggesting effective generalization abilities beyond real-world demographics; (3) Instruction-following capability gradually declines as dialogue turns increase, aligning with observations from previous studies Kwan et al.
 (2024); Maharana et al.
@@
 Appendix D Prompt Templates
-Case Study on Authenticity of Simulated Dialogue by Different User Simulators ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): Hey there!
+Human Evaluation Guidelines for Authenticity 1.
+Task Description: Please choose which user in the two test conversations is more similar to the reference conversation being spoken by the same person.
+2.
+Evaluation Criteria: Semantic Similarity: Measure the thematic consistency and discourse coherence between the generated user utterance and the target user utterance.
+Preference should be given to the utterance that more accurately reflects the semantic content of the target.
+Stylistic Parity: Analyze whether the generated user utterance matches the style of the target user utterance, including its tone, vocabulary, and grammatical structure.
+The utterance that aligns more closely with the stylistic elements of the target should be favored.
+Quality: Examine the fluency and logical coherence of the user utterance, focusing on the linguistic and logical smoothness of the user utterance.
+The more coherent and fluent utterances should be chosen.
+Figure 7: Human evaluation guidelines for authenticity.
+Human Evaluation Guidelines for Consistency 1.
+Task Description: Select the dialogue that contains the most appropriate user utterance from the two given generated dialogues based on the target user profile.
+2.
+Evaluation Criteria: Persona Reflection: How well does the utterance capture the profile’s themes, tone, and style?
+Prefer utterances that align closely with the profile’s characteristics.
+Comprehensiveness: Does the utterance integrate both objective facts and subjective traits?
+Favor broader, more complete reflections over narrow ones.
+Quality: Evaluate fluency, coherence, and naturalness.
+Prefer utterances that read smoothly, are logically consistent, and resemble genuine human conversation.
+Figure 8: Human evaluation guidelines for consistency.
+Case Study on Authenticity of Simulated Dialogue by Different User Simulators GPT-4o (User w/ Profile) Simulated Dialogue: GPT-4o (User w/ Profile): Hey there!
 I’m really excited to dive into some AI discussions with you… Assistant: Hello!
-I’m thrilled to chat about AI with you… ProfileGPT(4o): It’s great to dive into these with someone who shares such a passion… Assistant: I’m glad to hear that you’re as passionate about AI as I am!
-Your interests… ProfileGPT(4o): I’d be delighted to guide you in translating these advanced AI concepts… Assistant: I’m excited to help guide your ideas into actionable solutions … ProfileGPT(4o): I’m thrilled at the prospect of collaborating to develop solutions… Assistant: I’m excited at the prospect of working together to make these AI strategies… PlatoLM Simulated Dialogue: PlatoLM: how to train a model like you Assistant: I am a large language model… preprocessing step involves tokenizing the text… PlatoLM: what do you mean by tasks such as tokenizing text… explain these terms please?
+I’m thrilled to chat about AI with you… GPT-4o (User w/ Profile): It’s great to dive into these with someone who shares such a passion… Assistant: I’m glad to hear that you’re as passionate about AI as I am!
+Your interests… GPT-4o (User w/ Profile): I’d be delighted to guide you in translating these advanced AI concepts… Assistant: I’m excited to help guide your ideas into actionable solutions … GPT-4o (User w/ Profile): I’m thrilled at the prospect of collaborating to develop solutions… Assistant: I’m excited at the prospect of working together to make these AI strategies… PlatoLM Simulated Dialogue: PlatoLM: how to train a model like you Assistant: I am a large language model… preprocessing step involves tokenizing the text… PlatoLM: what do you mean by tasks such as tokenizing text… explain these terms please?
 Assistant: Certainly!
@@
 Assistant: Training and deploying your own AI model involves several key steps… Reference Dialogue: User: how to train a model like you Assistant: I am a large language model, which means that I was trained on… User: what is your base model Assistant: I am a large language model called GPT… User: isn’t you vicuna Assistant: I am not aware of any information or research about a model called "vicuna"… User: what hardware do i need to run you Assistant: I am a large language model called GPT, which was trained on…
-Figure 10: Case study comparing USP with other user simulators in the first four turns (out of 10 turns).
-USP and ProfileGPT(4o) depend exclusively on the provided profile, while PlatoLM utilizes the first-turn context.
-All simulators interact with GPT-4o for dialogue generation, and the reference dialogue below serves as the target for reconstruction.
+Figure 9: Case study comparing USP with other user simulators over the first four of ten dialogue turns.
+USP and GPT-4o (User w/ Profile) rely solely on the given profile, while PlatoLM uses the first-turn golden context.
+All simulators interact with GPT-4o, aiming to reconstruct the reference dialogue shown below.
 Case Study on Consistency of Simulated Dialogues across Different User Simulators Profile: You are engaging in a conversation with an AI assistant.
@@
 You often find yourself thinking of five-letter words with tricky placements, such as words containing ’A’ in the fourth position, or concocting ones with ’R’ in the second spot and ’T’ at the end.
-This aligns with your appreciation for linguistic puzzles and an enduring enjoyment of wordplay, reflecting your cognitive curiosity and dedication to enhancing your language skills.
-Your personality shines through a playful and engaging language style, particularly evident when you use structured questions to create dynamic game scenarios.
-You approach life with a vivid imagination, delighting in intellectual challenges that require intricate problem-solving.
-Your attention to detail is noteworthy, and you thrive in situations where organization is key.
-Preferring a more reserved lifestyle, you cherish moments of solitude and introspection, embracing activities that allow you to delve into complex, solitary intellectual pursuits with enthusiasm and precision.
+This aligns with your appreciation for linguistic puzzles and an enduring enjoyment of wordplay, reflecting your cognitive curiosity to enhancing your language skills.
+Your vibrant personality shines through a playful, engaging language style, crafting dynamic game scenarios with structured questions.
+You tackle intellectual challenges with vivid imagination and meticulous problem-solving.
+Highly organized and detail-oriented, you excel in complex tasks.
+Embracing a reserved lifestyle, you find joy in solitary, introspective pursuits, diving into intricate intellectual endeavors with enthusiasm and precision.
 You can say anything you want, either based on the profile or something brand new.
-ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): How about a playful journey into the world of word puzzles?
+GPT-4o (User w/ Profile) Simulated Dialogue: GPT-4o (User w/ Profile): How about a playful journey into the world of word puzzles?
 Let’s start with… Assistant: Sure!
@@
 The letter ’A’ is in the fourth position.
-ProfileGPT(4o): Wonderful!
-"Brave" is both intriguing and inspiring.
+GPT-4o (User w/ Profile): Wonderful!
+"Brave" is intriguing.
 Now, let’s spice it up… Assistant: Absolutely!
-The word "crest" fits your criteria, with ’R’… ProfileGPT(4o): Splendid!
+The word "crest" fits your criteria, with ’R’… GPT-4o (User w/ Profile): Splendid!
 "Crest" evokes such a vivid image; I love it….
 Assistant: Of course!
-The word "shine" fits your criteria, with ’S’ as the first letter … ProfileGPT(4o): Brilliant choice!
+The word "shine" fits your criteria, with ’S’ as the first letter … GPT-4o (User w/ Profile): Brilliant choice!
 "Shine" has such a radiant quality to it… Assistant: I’m glad you’re enjoying it!
@@
 Assistant: "GNW" is a four-letter word.
-Let’s think of another five-letter word with "A" in the… USP: Next Question: Give me one example of a 6-letter word having R at second place and T at sixth place.
+Let’s think of another five-letter word with "A" in the… USP: Next Question: Give me an example of a 6-letter word with R at second place and T at sixth place.
 Assistant: Sure!
 A six-letter word with "R" in the second position and "T" at the sixth position… USP: Great!
-Now give me one example of a 7-letter word having S at third place and H at seventh place.
+Now give me one example of a 7-letter word with S at 3rd place and H at 7th place.
 Assistant: Sure!
@@
 Answer: Assistant: Panda.
-User: Question: Suggest a five letter word which has R in second position and T in fifth position.
+User: Question: Suggest a 5 letter word with R in second position and T in fifth position.
 Answer: Assistant: Graft.
-Figure 11: Case study of user simulator comparisons showing the first four turns of a 10-turn dialogue.
-USP and other simulators interact with GPT-4o using only the provided profile, with the reference dialogue serving as the reconstruction target.
-Prompt for Extracting Big Five Personality Traits Task: Analyze the provided dialogue to assess the user’s personality traits across 5 personality dimensions.
-Focus exclusively on the user’s characteristics, disregarding any information related to others individuals, unless it directly impacts the user.
-For each personality trait: 1.
-Assign each dimension independently with score: "High", "Low," or "Inconclusive." 2.
-Provide conclusion: A high-level description, with concise supporting details.
-3.
-Provide reason: Justify the assigned score with specific evidence from the dialogue.
-4.
-Mark traits as Inconclusive only when no clear evidence exists.
-Personality Trait Defination: {{metric}}: {{definition}} Format:
-    "Trait Name": {
-        "score": "High/Low/Inconclusive",
-        "conclusion": "The user is a [trait descriptor] person...",
-        "reason": "Explanation referencing specific dialogue evidence."
-Example:
-[User]: "She is my age, in a homeless women’s shelter, living under very poor
-conditions. She is a mental health client, but the treatment team seems to
-... Her background is similar to mine, and I cannot abandon her."
-Detected Personality Traits:
-  "Conscientiousness": {
-    "score": "High",
-    "conclusion": "The user is a conscientious person who demonstrates a
-    sense of duty and commitment.",
-    "reason": "The user expresses a strong sense of responsibility ..."
-  "Agreeableness": {
-    "score": "High",
-    "conclusion": "The user is an empathetic and compassionate person who
-    values relationships.",
-    "reason": "The user shows care and concern for their cousin’s well-being..."
-  "Extraversion": {
-    "score": "Inconclusive"
-Figure 12: Prompt for extracting deep intrinsic characteristics.
+Figure 10: Case study comparing user simulators over the first four turns of a 10-turn dialogue.
+USP and other simulators interact with GPT-4o using only the provided profile, targeting the reference dialogue for reconstruction.
 Prompt for Extracting Scene-Consistent Attributes User Persona Analysis Task Objective The primary goal of this task is to analyze user utterances in-depth and accurately extract key persona attributes based on both direct and implicit cues.
@@
 Each task should be concisely summarized with specifics.
-If there are multiple tasks or requests, list each one separately.
+If there are multiple tasks, list each separately.
 Example User Utterances:
-[User]: Summarize: Harry Potter is a fictional character in the Harry Potter series...
+[User]: Summarize: Harry Potter is a fictional character in Harry Potter series...
 [Assitant]: Harry Potter is a fictional character...
@@
 Figure 12: Prompt for extracting scene-related attributes.
+Prompt for Extracting Big Five Personality Traits Task: Analyze the provided dialogue to assess the user’s personality traits across 5 personality dimensions.
+Focus exclusively on the user’s characteristics, disregarding any information related to others individuals, unless it directly impacts the user.
+For each personality trait: 1.
+Assign each dimension independently with score: "High", "Low," or "Inconclusive." 2.
+Provide conclusion: A high-level description, with concise supporting details.
+3.
+Provide reason: Justify the assigned score with specific evidence from the dialogue.
+4.
+Mark traits as Inconclusive only when no clear evidence exists.
+Personality Trait Defination: {{metric}}: {{definition}} Format:
+    "Trait Name": {
+        "score": "High/Low/Inconclusive",
+        "conclusion": "The user is a [trait descriptor] person...",
+        "reason": "Explanation referencing specific dialogue evidence."
+Example:
+[User]: "She is my age, in a homeless women’s shelter, living under very poor
+conditions. She is a mental health client, but the treatment team seems to
+... Her background is similar to mine, and I cannot abandon her."
+Detected Personality Traits:
+  "Conscientiousness": {
+    "score": "High",
+    "conclusion": "The user is a conscientious person who demonstrates a
+    sense of duty and commitment.",
+    "reason": "The user expresses a strong sense of responsibility ..."
+  "Agreeableness": {
+    "score": "High",
+    "conclusion": "The user is an empathetic and compassionate person who
+    values relationships.",
+    "reason": "The user shows care and concern for their cousin’s well-being..."
+  "Extraversion": {
+    "score": "Inconclusive"
+Figure 13: Prompt for extracting deep intrinsic characteristics.
 Prompt for Rephrasing Attributes into Natural Descriptions Narrative Generation Objective Rephrase the provided key-value pairs into a natural, coherent second-person description.
@@
 Figure 14: Prompt for rephrasing attributes into natural descriptions for profile generation.
-Human Evaluation Guidelines for Authenticity 1.
-Task Description: Please choose which user in the two test conversations is more similar to the reference conversation being spoken by the same person.
-2.
-Evaluation Criteria: Semantic Similarity: Measure the thematic consistency and discourse coherence between the generated user utterance and the target user utterance.
-Preference should be given to the utterance that more accurately reflects the semantic content of the target.
-Stylistic Parity: Analyze whether the generated user utterance matches the style of the target user utterance, including its tone, vocabulary, and grammatical structure.
-The utterance that aligns more closely with the stylistic elements of the target should be favored.
-Quality: Examine the fluency and logical coherence of the user utterance, focusing on the linguistic and logical smoothness of the user utterance.
-The more coherent and fluent utterances should be chosen.
-Figure 16: Human evaluation guidelines for authenticity.
-Human Evaluation Guidelines for Consistency 1.
-Task Description: Select the dialogue that contains the most appropriate user utterance from the two given generated dialogues based on the target user profile.
-2.
-Evaluation Criteria: Persona Reflection: Evaluate how closely each user utterance reflects the target profile’s thematic content, tone, and stylistic features.
-Preference should be given to the dialogue whose user utterance most accurately reflects the profile’s characteristics in terms of thematic coherence and expression.
-Comprehensiveness: Assess the extent to which the user utterance encapsulates the target profile, integrating both objective facts and subjective descriptions.
-The more comprehensive utterance, which addresses a broader range of profile aspects, should be favored over one that focuses narrowly on a single dimension.
-Quality: Evaluate the fluency, coherence, and human-likeness of the user utterance.
-Preference should be given to the utterance that demonstrates greater linguistic smoothness, logical cohesion, and alignment with genuine human conversational patterns.
-Figure 17: Human evaluation guidelines for consistency.
-Prompt for NLI Score of Profile Precision Based on Given Dialogue Role You are an expert in evaluating the consistency between a given user profile (Source) and the user’s utterance (Target).
+Prompt for NLI Score of Profile Precision(DP.P) Based on Given Dialogue Role You are an expert in evaluating the consistency between a given user profile (Source) and the user’s utterance (Target).
 Your task is to assess whether the Target aligns with, contradicts, or is ambiguous in relation to the Source.
@@
 Concise & Precise Justification: The reasoning should be clear, objective, and free from unnecessary elaboration.
-Figure 18: Prompt for DP.P based on NLI.
-Prompt for NLI Score of Dialogue Precision Based on Given Profile Role You are an expert in evaluating consistency between a given dialogue history (Source) and a corresponding user profile description (Target).
+Figure 15: Prompt used by GPT-4o for NLI-based evaluation of DP.P and r-DP.R metrics.
+Prompt for NLI Score of Dialogue Precision(r-DP.P) Based on Given Profile Role You are an expert in evaluating consistency between a given dialogue history (Source) and a corresponding user profile description (Target).
 Your task is to determine whether the Target aligns with, contradicts, or is ambiguous in relation to the Source.
@@
 Concise & Precise Justification: The reasoning should be clear, objective, and free from unnecessary elaboration.
-Figure 19: Prompt for DP.R based on NLI.
+Figure 16: Prompt used by GPT-4o for NLI-based evaluation of DP.R and r-DP.P metrics.
 Prompt for Subjective Characteristics Score (SC.Score) in Consistency Evaluation Task Description You are tasked with evaluating the quality of user responses in real human-LLM interactions.